### v1

# arXiv:2603.01316v1[eess.AS] 01 Mar 2026

Abstract
This paper investigates the use of relative cues for text-based target speech extraction (TSE). We first provide a theoretical justification for relative cues from the perspectives of human perception and label quantization, showing that relative cues preserve fine-grained distinctions often lost in absolute categorical representations. Building on this analysis, we propose a two-stage TSE framework, in which a speech separation model generates candidate sources, followed by a text-guided classifier that selects the target speaker based on embedding similarity. Using this framework, we train two separate classification models to evaluate the advantages of relative cues over independent cues in terms of both classification accuracy and TSE performance. Experimental results demonstrate that (i) relative cues achieve higher overall classification accuracy and improved TSE performance compared with independent cues, (ii) the two-stage framework substantially outperforms single-stage text-conditioned extraction methods on both signal-level and objective perceptual metrics, and (iii) certain relative cues (language, gender, loudness, distance, temporal order, speaking duration, random cue and all cue) can surpass the performance of an audio-based TSE system. Further analysis reveals notable differences in discriminative power across cue types, providing insights into the effectiveness of different relative cues for TSE.
I Introduction
TSE systems leverage auxiliary information—such as speaker embeddings derived from an enrollment speech sample, spatial location, visual cues from lip movements, and textual descriptions of the target speaker’s attributes—to enable the TSE model to selectively attend to and extract the desired voice in a mixture of multiple speakers [65, 61]. Applications such as smart assistants, hearing aids, and meeting transcription benefit from TSE’s ability to isolate the desired voice, enhancing both accuracy and usability by leveraging speaker-specific cues [65, 61, 44, 2].
Recent years have witnessed rapid advances in target speech extraction, driven by progress in deep learning, self-supervised pretraining, and multimodal modeling. Early works in speaker-dependent extraction exploited the idea of conditioning the separation network on speaker identity, giving rise to architectures such as SpeakerBeam [64] and VoiceFilter [56]. These models employ a reference-guided framework, where a short clean utterance of the target speaker provides a speaker embedding that modulates the separation network to enhance the target voice.
Subsequent developments incorporated pretrained speaker encoders, such as D-Vectors[55], X-Vectors[51], and ECAPA-TDNN[8], to obtain robust speaker embeddings, decoupling speaker modeling from the extraction network. Moreover, multi-modal approaches integrating audio, video cues have gained attention[37, 26, 54, 58], especially for audiovisual datasets like AVSpeech[11] or VoxCeleb[35]. These methods align visual speech movements or lip information with audio stream[3, 25], improving robustness in visually grounded TSE scenarios.
Despite these advancements, several challenges remain, including robustness to unseen speakers, limited availability of clean target references in real-world scenarios, and mismatch between the enrollment speech and the mixture input regarding recording conditions, speaking style, or acoustic environment. Beyond acoustic/visual-based guidance, recent research has explored text-guided extraction. These models use a text query (prompt) to describe speech attributes of target speaker as a condition to guide the extraction process of the desired speaker’s speech signal in the mixture, enabling extraction when no enrollment utterance of target speaker is available. Compared to other cues, text offers greater accessibility, enabling users to provide instructions without recording audio or video. It also provides greater flexibility, as it can convey a wide range of target speaker characteristics, including linguistic cues (e.g., speech transcription), paralinguistic cues (e.g., gender), and acoustic cues (e.g., loudness of speech). In addition, text offers stronger privacy by avoiding the transmission or storage of personal voice data.
Hao et al. were the first to successfully introduce text-based TSE that used a LLM for extracting semantic information from text queries and integrating it into a speech separation network [16]. Subsequently, Huo et al. proposed a flexible text-guided TSE framework and introduced the TextrolMix dataset, which provides rich textual descriptions of speech attributes, including speaker identity, emotion, pitch level, gender, accent, and speaking rate [17]. Similarly, Seki et al. developed a TSE model conditioned on text prompts describing paralinguistic and non-linguistic attributes such as gender, pitch, speaking speed, and loudness, demonstrating that it can overcome the limitations of existing audio-based TSE methods [46].
Previous text-based TSE studies typically represent continuous-valued speech attributes, such as speaking rate, using a small number of discrete categories (e.g., slow, normal, fast) without considering inter-speaker relations. This is usually done by first estimating each speaker’s attribute values in a mixture and then quantizing them into predefined categories. Such coarse quantization can obscure meaningful inter-speaker differences, as within-category variation is discarded, reducing discriminative information. Consequently, speaker-independent cues lose finer-grained distinguishing information when the target and interfering speakers have subtle attribute differences but are assigned to the same category.
To overcome this limitation, our prior work introduced inter-speaker relative cues, which explicitly describe differences between the target and interfering speakers [7]. By formulating attributes relatively rather than as speaker-specific categories, these cues preserve fine-grained distinctions that are critical for target speaker extraction. To study these cues, we constructed a two-speaker mixture dataset spanning a wide range of attributes, including pitch level, pitch range, speaking rate, speaking duration, temporal order, age, loudness, speaker-to-microphone distance, language, gender, transcription and emotion [7].
This approach is motivated by human perception: people perceive continuous-valued speech attributes primarily through comparative listening, rather than by measuring precise physical quantities. Instead of relying on exact values—such as fundamental frequency ($F_{0}$) in Hertz or speaking rate in syllables per minute—listeners judge whether one speaker sounds higher or lower in pitch, or faster or slower in speaking rate than another [53, 33, 32]. Conventional TSE systems that rely on absolute, discretized labels fail to capture these relative distinctions. Subtle but perceptually meaningful differences between speakers can be lost when multiple speakers fall into the same category (e.g., both labeled as “fast” speaking rate). By explicitly encoding inter-speaker comparisons—such as “Speaker A has a faster speaking rate than Speaker B”—relative attribute cues preserve these contrasts, providing more discriminative and perceptually aligned conditioning for target speaker extraction.
Building upon these insights, the main contributions of this work are summarized as follows:
  • Theoretical Justification of Relative Cues: We provide a theoretical analysis of relative cues, which accounts for human perception and label quantization, offering a principled explanation of their advantages over speaker independent cues, which we hereafter refer to as independent cues throughout the remainder of this paper.
  • Proposed Two-Stage Text-Guided TSE Method: In our previous text-based TSE system—similar to conventional audio-based approaches—the conditioning-and-extraction paradigm is prone to speaker confusion [62, 28, 34, 10] during inference, leading to degraded performance. Motivated by enrollment-audio-based two-stage TSE approach [10], and to improve upon our previous single-stage system, we adopt a two-stage framework in this work: first, performing speech separation, and then using a classification model to select the target speaker based on inter-speaker relative cues embedded in text prompts. This approach not only mitigates the speaker confusion inherent in single-stage TSE systems but also provides interpretability, allowing insight into the relative effectiveness of different cues through analysis of the classification accuracy for each cue.
  • Empirical Evaluation of the Advantages of Relative Cues: Based on the proposed two-stage framework, we evaluate the impact of using relative versus independent cues on overall target speech classification and the corresponding TSE performance, focusing on continuous-valued attributes. This provides an empirical perspective on the advantages of relative cues over independent cues.
The remainder of this paper is organized as follows. Section II introduces the inter-speaker relative cues and presents their theoretical justification. Section III describes the proposed two-stage method. Section IV presents the dataset, training configuration, evaluation setup, experimental results, and corresponding analysis. Finally, Section V concludes the paper.
II Inter-Speaker Relative Cues
The proposed text-guided TSE system takes a two-speaker mixture as input along with a text prompt which consists of cues indicating the target speaker, aiming to extract the target speech signal. In this work, we employ inter-speaker relative cues, which describe the relative differences in speech attributes between the target speaker and the interfering speaker in the mixture. According to the value type of speech attributes, they can be categorized into discrete-valued and continuous-valued types.
Continuous-valued speech attributes correspond to estimations of specific speech or acoustic properties. In this study, these attributes include:
  1. 1.
    mean $\bm{F_{0}}$, computed as the average fundamental frequency over all voiced frames in an utterance;
  2. 2.
    $\bm{F_{0}}$ span, defined as the difference between the maximum and minimum $F_{0}$} within an utterance;
  3. 3.
    speaker age, treated as a continuous variable in years;
  4. 4.
    speaking duration, defined as the total active speech time (in seconds) of a speaker within the mixture;
  5. 5.
    speaking rate, measured as the number of syllables per minute in the speech signal;
  6. 6.
    RMS energy (dB), computed as $20\log_{10}$ of the root-mean-square amplitude of the speech signal;
  7. 7.
    speaker-to-microphone distance, expressed in meters; and
  8. 8.
    appearance time, defined as the onset time of a speaker’s speech within the mixture.
Discrete-valued speech attributes correspond to categorical properties that take values from a finite set. In this study, these attributes include:
  1. 1.
    language, indicating the spoken language of the utterance;
  2. 2.
    gender, representing the speaker’s perceived acoustical gender category;
  3. 3.
    emotional state, describing the affective category expressed in the speech signal; and
  4. 4.
    transcription content, corresponding to the linguistic content of the utterance in textual form.
We derive each inter-speaker relative cues by comparing a single speech attribute between the target speaker and the interfering speaker in a two-speaker mixture. Each relative cue encodes the difference along one specific attribute only. Independent cues, in contrast to relative cues, are derived solely from the attributes of an individual speaker—either the target or the interfering speaker—without accounting for inter-speaker relationships.
II-A Independent Cues
The key difference between relative and independent cues lies in how continuous-valued attributes are represented. Let $x_{i}^{(\mathrm{tar})}$ and $x_{i}^{(\mathrm{inf})}$ denote the value of a continuous-valued speech attribute $i$ (e.g., speaking rate) for the target and interfering speakers, respectively. Independent cues represent each continuous-valued attribute using a small number of discrete categories. The attribute space is segmented into these categories based on the distribution of attribute values across all individual target and interference speakers in the training set. Each attribute value $x_{i}$ is quantized as
$\tilde{x}_{i}=\mathcal{Q}_{i}(x_{i}),\quad\tilde{x}_{i}\in\{\mathit{A}_{i,1},\mathit{A}_{i,2},\dots,\mathit{A}_{i,k_{i}}\},$ (1)
where $\mathcal{Q}_{i}(\cdot)$ denotes the quantization operator for the $i$-th attribute, $k_{i}$ is the number of discrete categories for that attribute, and $\mathit{A}_{i,1},\dots,\mathit{A}_{i,k_{i}}$ represent the discrete categories (e.g., low, normal, and high for mean $F_{0}$).
The corresponding independent cue category for attribute $i$ is given by $\tilde{x}_{i}$. This mapping from continuous-valued attribute values $x_{i}$ to discrete independent cue categories $\tilde{x}_{i}$ is illustrated in Fig. 1(a). Due to the coarse nature of the categories, perceptually noticeable differences in the original continuous-valued attribute may be mapped to the same discrete category (as illustrated by the shaded region in Fig. 1(a)), in which case the attributes of two speakers cannot be distinguished from each other.
(a) Independent Cue
(b) Relative Cue for Absolute Difference
(c) Relative Cue for Percentage Difference
Figure 1: Illustration of quantization schemes for continuous-valued attributes. (a) Independent cues: the attribute value space of the target and interfering speakers, denoted by $x_{i}^{(\mathit{tar})}$ (y-axis) and $x_{i}^{(\mathit{inf})}$ (x-axis), is partitioned into a small number of coarse categories (three shown as in the example). Shaded regions indicate attribute values of the two speakers that are quantized into the same category, in which case the two speakers cannot be distinguished based on the attribute. (b) Relative cues for absolute differences: attribute values of the target and interfering speakers are compared with absolute difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Regions outside the shaded area represent perceptually distinct absolute differences. (c) Relative cues for percentage differences: attribute values of the target and interfering speakers are compared with percentage difference. The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold. Because the comparison is based on relative differences, the Similar region gradually expands from the zero point. Regions outside the shaded area represent perceptually distinct relative differences.
II-B Relative Cues
In contrast, relative cues categorize continuous-valued attributes into groups (e.g., slower, similar, and faster for speaking rate) by comparing the attribute values of the two speakers. For the attributes RMS energy (dB), speaker-to-microphone distance, speaker age, and appearance time in the mixture, the difference is expressed as the absolute difference
$\Delta x_{i}^{|.|}=x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}.$ (2)
These attributes correspond to interpretable physical or temporal quantities whose perceptual relevance is largely determined by their absolute values. Changes in these quantities can be directly perceived and interpreted by human listeners without reference to a baseline. For attributes such as mean $F_{0}$, $F_{0}$ span, speaking rate, and speaking duration, we employ the percentage difference
$\Delta x_{i}^{\%}=\frac{x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}}{\min\!\left(x_{i}^{(\mathrm{tar})},x_{i}^{(\mathrm{inf})}\right)}\times 100\%.$ (3)
These attributes represent global prosodic properties, for which human listeners are more sensitive to proportional than absolute changes [12, 23, 52], making this formulation better aligned with human auditory perception across different baselines.
For notational convenience, we introduce the variable $\Delta x_{i}^{(*)}\in\{\Delta x_{i}^{|.|},\Delta x_{i}^{\%}\}$ to denote the attribute-specific difference measure. This measure is then compared against the predefined threshold $\theta_{i}$ of each attribute, which are shown in Table I.
TABLE I: Continuous-Valued Speech Attributes with Our Defined Thresholds for Relative Cues and Motivated Speech Perceptual Studies
Speech Attribute Threshold ($\theta_{i}$) Motivated Findings
RMS energy (dB) 3 dB Perception of relative loudness differences in speech is mediated by changes in loudness-related acoustic features, such as excitation strength derived from the glottal source [47]. McShefferty et al. [30] report that the average just-noticeable difference (JND) in speech-to-noise ratio is approximately 3 dB. Signal-to-interference ratio (SIR), computed based on RMS energy over active speech regions, can serve as an approximation of the perceived loudness difference between target and interfering speech.
Speaker-to-microphone distance 0.5 m Human perception of speaker-to-microphone distance relies on relative variations in distance-dependent acoustic cues, particularly changes in the direct-to-reverberant energy ratio [48, 31, 59].
Speaker age 10 years Perception of speaker age is based on relative differences in age-related vocal characteristics, such as vocal tract resonance and speaking rate. Listeners typically judge speakers as “younger” or “older” rather than estimating absolute age [27, 50, 57].
Mean $F_{0}$ 6% Pitch perception is governed by relative changes in F0. Listeners’ sensitivity to differences in F0 changes in speech and found that for many listeners, average JNDs were around 1.5–2 semitones (6-12%) [1].
$F_{0}$ span 25% Perception of pitch range ($F_{0}$ span) depends on relative variation in pitch-related features. Wider pitch ranges are often associated with high arousal or positive affect, while narrower ranges indicate low arousal [19, 5, 45, 20].
Speaking rate 15% Humans perceive differences in speaking rate through relative tempo changes, with a JND of approximately 5% of the base rate [40]. Descriptors such as “much slower” or “faster” reflect this relative perception [13, 41].
Speaking duration 15% Perception of speech duration follows Weber’s law, whereby the JND scales proportionally with the original duration, indicating reliance on relative temporal differences [15].
Appearance time 0.1 s
Some thresholds are motivated by directly adopting suggested JNDs or by adjusting them to larger values based on findings from prior psychoacoustic and speech perception studies. The remaining thresholds, which either lack suggested JNDs or have no related studies, are defined by us [^1] These thresholds are largely chosen in an ad hoc manner, as no perceptual studies have used exactly the same experimental setup..
The relative cue category $r_{i}$ for attribute $i$ is assigned as
$r_{i}=\begin{cases}\mathit{R}_{i,1},&\Delta x_{i}^{(*)}>\theta_{i},\\[2.0pt] \mathit{Similar},&|\Delta x_{i}^{(*)}|\leq\theta_{i},\\[2.0pt] \mathit{R}_{i,2},&\Delta x_{i}^{(*)}<-\theta_{i},\end{cases}$ (4)
where $\mathit{R}_{i,1}$, $\mathit{Similar}$, and $\mathit{R}_{i,2}$ denote relational categories corresponding, for example, to longer, similar, and shorter speaking duration relative to the interfering speaker. This mapping is illustrated in Fig. 1(b) and Fig. 1(c). The distribution of these relative cues under different threshold settings in the test set (based on the simulated dataset described in Section IV) are presented in Table II. The total number of samples is almost balanced across cue types, except for speaker age, for which labeled samples are limited in our dataset.
TABLE II: Sample counts of relative categories for each continuous attribute in the test set, based on the defined thresholds.
Continue-valued Attribute R1 Similar R2
RMS energy 2547 4992 2461
Speaker-to-microphone distance 1441 7124 1435
Speaker age 198 605 197
Mean $F_{0}$ 4394 1081 4515
$F_{0}$ span 3799 2376 3813
Speaking rate 3331 2235 3465
Speaking duration 3881 2227 3892
Appearance time 3179 3511 3310
II-C Discrete-Valued Attributes
For discrete-valued attributes, independent cues specify the attribute value (category) of the target speaker (e.g., the language spoken) and do not consider whether it is shared with the interfering speaker. In contrast, relative cues indicate whether the target and interfering speakers share the same attribute value (category) or are distinct with respect to that attribute. Let $d_{i}^{(\mathrm{tar})}$ and $d_{i}^{(\mathrm{inf})}$ denote the values of discrete-valued attribute $d_{i}$ for the target and interfering speakers, respectively, with $i$ indexing a particular attribute (e.g., language). Then the independent cue category $a_{i}$ for the target speaker is defined as
$a_{i}=d_{i}^{(\mathrm{tar})},$
while the relative cue category $r_{i}$ for the target speaker is defined as
$r_{i}=\begin{cases}\text{Same},&d_{i}^{(\mathrm{tar})}=d_{i}^{(\mathrm{inf})},\\[2.84526pt] d_{i}^{(\mathrm{tar})},&d_{i}^{(\mathrm{tar})}\neq d_{i}^{(\mathrm{inf})},\end{cases}$
so that when the attribute values (categories) differ, the relative cue retains the original categorical distinction of the target speaker.
II-D Text Prompt Generation
Inter-speaker relative cues are formulated as natural-language descriptions (text prompts) that explicitly characterize the relative differences between the target and interfering speakers. These prompts are used for both training and evaluation of the TSE systems. Each text prompt is generated by populating predefined templates with one or more relative cue descriptors, following the same approach as [7]. The templates incorporate action-oriented verbs such as extract, separate, or isolate, combined with descriptive phrases that emphasize the distinguishing characteristics of the target speaker. For example, a cue vector indicating (female, higher pitch, faster speaking rate) can be verbalized as: “Please extract the female speaker with a higher pitch and faster speaking rate.” We also employed the GPT-4o-mini model to automatically rephrase the templates into more generic prompts. Although these diversified prompts increased linguistic variability, they resulted in lower performance. Therefore, we adopt the original template-based prompts in our experiments. Since relative cues are mixture-specific, the target speaker in a given mixture may be specified by one or multiple relative cues.
Table IV reports the number of mixtures in each subset that contain a given relative cue. Since a single mixture may include multiple cues, the counts for individual cues do not sum to the total number of mixtures. As shown in the table, the target speaker in each mixture can be specified using different types of prompts. The Random Cue type consists of a combination of randomly selected relative cues, where the number of selected cues ranges from 2 to (n1)(n-1) and $n$ denotes the total number of relative cues in the complete set. In contrast, the All Cue type includes all $n$ relative cues for a given mixture. This procedure yields a diverse set of natural-language descriptions that accurately reflect the underlying attribute differences between speakers.
III Proposed Two-Stage Method
Figure 2: Overview of the proposed two-stage text-guided TSE system during inference.
In this work, we adopt a two-stage approach for text-guided target speech extraction, comprising: (1) speech separation and (2) target speech classification guided by a text prompt. This design is inspired by prior two-stage target speech extraction framework, which first separate mixture speech signals that perform speech separation followed by speaker verification based on speaker embeddings derived from enrollment utterances [10]. Such a cascaded design [10] has been shown to improve TSE performance compared to single-stage TSE systems. In this two-stage framework, the two components are trained independently, which simplifies the optimization process and allows each module to focus on a more specific objective. This modular strategy leads to a more stable and interpretable architecture than single-stage TSE approaches. An overview of the system during inference is provided in Fig. 2: the input mixture is first separated into individual speech signals, after which a text-guided target selection mechanism identifies and outputs the desired speech signal.
Let the input mixture be denoted as mixmix, which contains two speakers. Given a text prompt $p$ describing one or more relative cues of the target speaker, the goal is to extract the target speech signal from mixmix.
Stage 1: Speech Separation
In the first stage, the mixture is separated into individual speech signals using a speech separation model $\mathcal{S}(\cdot)$
$\hat{s}_{1},\hat{s}_{2}]=\mathcal{S}(\text{mix}),$
trained with utterance-level permutation invariant training (PIT) [22], where $\hat{s}_{1}$ and $\hat{s}_{2}$ denote the estimated speech signals of the individual speakers. PIT resolves the permutation ambiguity by computing the loss for the two possible output-reference assignments and selecting the minimum. This allows the model to learn speaker-specific separation without assuming a fixed output order. The separation stage produces isolated speech signals that serve as input for the subsequent target speech classification stage.
Stage 2: Target Speech Classification via Text Prompt
The second stage identifies which separated signal corresponds to the target speaker using the provided text prompt.
Speaker and Text Embeddings
Each separated signal $\hat{s}_{j}$ (where $j\in\{1,2\}$) is encoded by an audio encoder to produce a fixed-dimensional embedding as
$\tilde{z}_{j}=\mathcal{A}(\hat{s}_{j}).$
The audio encoder consists of a self-supervised speech model (wav2vec2-large-xlsr-53222https://huggingface.co/facebook/wav2vec2-large-xlsr-53), followed by a pooling operation, ReLU activation, and layer normalization.
Similarly, the text prompt $p$ is encoded into the same embedding space using a text encoder $\mathcal{T}$ as
$\tilde{z}_{p}=\mathcal{T}(p),$
which includes an LLM-based model (LLaMA 3.2 1B Instruct [^3] https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct) followed by a trainable linear projection with ReLU and layer normalization, mapping the LLM output to the speaker embedding space to align textual and speaker representations in a shared, directly comparable latent space.
Similarity Computation and Classification
The similarity of the text prompt and each separated signal is measured via the cosine similarity of their embeddings as
$\text{sim}_{j}=\frac{\tilde{z}_{p}\cdot\tilde{z}_{j}}{\|\tilde{z}_{p}\|\,\|\tilde{z}_{j}\|},\quad j\in\{1,2\}.$
We define the logit as the difference between these similarity scores as
$\text{logit}=\text{sim}_{1}-\text{sim}_{2},$
which represents how much more similar the first separated signal $\hat{s}_{1}$ is to the text prompt compared with the second signal $\hat{s}_{2}$. The logit is then mapped to a probability using the sigmoid function:
$\hat{y}=\sigma(\text{logit}),$
where $\hat{y}\in[0,1]$ indicates the likelihood that $\hat{s}_{1}$ corresponds to the target speaker described by the text prompt. Finally, the predicted target speaker index is determined as
$\text{pred}=\begin{cases}1,&\hat{y}>0.5,\\ 0,&\text{otherwise}.\end{cases}$
During inference, the signal with the higher similarity is selected as the target speech $\hat{s}_{\text{tar}}$.
Training Label Generation and Loss
To train the classifier, ground-truth labels are generated for each two-speaker mixture by comparing the separated signals with the reference target speech $s_{\mathrm{ref}}$ using scale-invariant signal-to-distortion ratio (SI-SDR) [24] as
[…]
These labels indicate which separated signal corresponds to the target speaker. The classifier is trained by minimizing the binary cross-entropy (BCE) loss between the predicted probability and the generated label:
$\mathcal{L}=\mathrm{BCE}(\hat{y},y),$
encouraging the model to correctly associate the text prompt with the separated speech signal of the target speaker.
IV Evaluation
The evaluation aims to: (i) assess the advantages of the proposed two-stage framework over conventional single-stage systems, and (ii) examine whether relative cues yield better performance than independent cues. The following sections provide an overview of the experimental dataset and training configurations, followed by detailed descriptions of each experiment, including the setup, results, and analysis.
IV-A Dataset
To ensure broad linguistic and acoustic diversity, audio mixtures were constructed from multiple publicly available corpora spanning five major languages—Chinese, English, French, German, and Spanish—each providing annotated attributes such as language, gender, emotion, age, and transcription (see Table II). Each mixture contains speech from two speakers—one designated as the target and the other as the interference—together with a text prompt describing their relative characteristics.
(1) Extraction of Speech Attributes: To provide ground-truth labels for the relative cue categories in each mixture, a set of speech attributes is extracted for every utterance. These attributes serve as the basis for constructing inter-speaker relative cues (e.g., higher pitch, faster speaking rate). All attributes are computed from the isolated clean speech signals prior to mixing, ensuring that the measurements are not affected by interference from other speakers or by mixture-level acoustic artifacts.
Mean $F_{0}$ and $F_{0}$ span are extracted using the Librosa pYIN library [^4] https://librosa.org/doc/main/generated/librosa.pyin.html. When computing speaking duration, word- and pause-level durations are first obtained. If transcription annotations are available, these durations are extracted from time boundaries provided by the Montreal Forced Aligner [^5] https://montreal-forced-aligner.readthedocs.io/en/latest/; otherwise, they are estimated using the Silero voice activity detection tool [^6] https://github.com/snakers4/silero-vad. Speaking duration is then defined as the sum of word durations and short pauses, excluding pauses longer than 0.6 s, which are treated as silence. Speaking rate is computed as the total number of syllables divided by the speaking duration. Syllable counts are estimated based on vowel sequences: in Chinese, each character is treated as one syllable, whereas in English, French, German, and Spanish, syllables are approximated by counting contiguous vowel groups within words. The RMS energy is computed from the reverberant speech while accounting for the speaking duration. The speaker-to-microphone distance is randomly sampled in a virtual room, with horizontal distances of 0.3–1.5 m to the microphone. Language, gender, age, emotional state, and transcription content are obtained directly from the original dataset annotations.
(2) Audio Mixtures: Each corpus is split into non-overlapping training, validation, and test subsets, with no overlap in speakers or spoken words. We generated 100,000 mixtures for training and 10,000 mixtures each for validation and testing. To create realistic speech mixtures, each mixture is composed of two speech signals, $S_{1}$ and $S_{2}$, from different speakers. Both signals are first processed to remove silence—using forced alignment for transcribed speech or voice activity detection for untranscribed segments—and then temporally arranged with partial overlaps to mimic conversational dynamics. If either $S_{1}$ or $S_{2}$ is shorter than 3 s, the overlap duration is set to the length of the shorter signal, and zero-padding is applied at the beginning of the shorter signal using a randomly selected start offset uniformly drawn from $0,\text{duration of longer sample}-\text{duration of shorter sample}]$. For longer signals, the overlap duration is set to $(\text{length}(S_{1})+\text{length}(S_{2})-6)$ s, with $S_{2}$ padded using a fixed start offset of $6-\text{length}(S_{2})$ s.
To incorporate realistic room acoustics, each $S_{1}$ and $S_{2}$ is convolved with a separate room impulse response (RIR) from the gpuRIR library [9], simulating different speaker positions within the room. The simulated rooms vary in size, reverberation time, and speaker–microphone distance. Room dimensions are randomly sampled within 9–11 m in length and width, and 2.6–3.5 m in height, while reverberation times (RT60) are sampled from 0.3–0.6 s. The microphone is positioned at the center of the room, and speaker locations are randomly assigned such that horizontal speaker–microphone distances range from 0.3 to 1.5 m, with speaker heights between 1.6 and 1.9 m. We keep the reverberant $S_{2}$ fixed and adjust the reverberant $S_{1}$ by applying a signal-to-interference ratio (SIR), uniformly sampled from –6 to 6 dB. For each mixture, one of the reverberant signals, $S_{1}$ or $S_{2}$, is randomly selected to serve as the target speech.
TABLE III: Summary of used speech datasets across five languages, detailing the number of speakers in the training, validation, and test splits, along with their partial attributes.
Dataset Language Speakers Attributes
CASIA [60] Chinese 2/1/1 Language, Emotion, Gender, Transcription
ESD [63] Chinese 6/2/2 Language, Emotion, Gender, Transcription
Aishell3 [49] Chinese 120/30/30 Language, Gender, Transcription
MagicData-Read [36] Chinese 33/8/11 Language, Age, Gender
Librispeech [38] English 120/30/30 Language, Gender, Transcription
ESD [63] English 6/2/2 Language, Emotion, Gender, Transcription
OreauFR-02 [21] French 21/7/4 Language, Emotion, Age, Gender
MLS French [39] French 90/18/18 Language, Gender, Transcription
EmoDB [6] German 6/2/2 Language, Emotion, Age, Gender, Transcription
MLS German [39] German 120/30/30 Language, Gender, Transcription
MLS Spanish [39] Spanish 79/20/20 Language, Gender, Transcription
EmoMatchSpanish [14] Spanish 30/10/10 Language, Emotion, Gender, Transcription
(3) Text Prompts: As detailed in Section II, Table IV summarizes the distribution of samples across relative cue categories in the training, validation, and test sets, excluding the “Similar” and “Same” category, as they do not offer discriminative power.
TABLE IV: Sample counts of each relative cue in the training, validation and test sets
Relative Cue Training Validation Test
Random Cue 99641 9962 9956
All Cue 99960 9996 9997
Language 73845 7192 7157
Transcription 92311 9186 9189
Gender 49114 5759 5125
Emotion 20377 1935 2012
Speaking Rate 67821 6674 6796
Speaking Duration 76498 7880 7773
Pitch Level 89366 8917 8909
Pitch Range 77019 7715 7612
Distance 27983 2672 2876
Age Category 3130 348 395
Loudness 50731 4971 5008
Temporal Order 60625 6686 6489
IV-B Training Configurations
For the speech encoder used in the classification model of the proposed two-stage framework, all cue types employed during training span linguistic, paralinguistic, and acoustic information. Therefore, the speech encoder plays a central role in capturing these diverse representations. We adopt the encoder from the pretrained wav2vec2-large-xlsr-53 model, a multilingual self-supervised speech representation model based on the wav2vec 2.0 architecture [4]. Trained on speech data from 53 languages, including Multilingual LibriSpeech and other large-scale corpora, wav2vec2-large-xlsr-53 learns universal, language-independent representations directly from raw audio. Compared to the state-of-the-art speaker encoder ECAPA-TDNN, which is primarily designed to extract speaker-specific acoustic embeddings, wav2vec2-large-xlsr-53 provides richer representations that jointly capture linguistic, paralinguistic, and acoustic information. This makes it particularly well suited for cross-lingual target speech extraction guided by diverse cue types. For wav2vec2-large-xlsr-53, only the first five encoder layers are used, with fine-tuning applied to the fifth layer.
To encode text prompts into embeddings, we employ the LLaMA 3.2 1B Instruct model for all single-stage TSE systems as well as for the classification model in the two-stage framework. The LLaMA 3.2 1B instruction model was fine-tuned using LoRA (rank 16, scaling factor 16, dropout 0.05) on the query and key projection layers of self-attention.
To ensure high-quality signal separation, we adopt the SOTA speech separation model TF-Locoforme [43], which has demonstrated superior performance on public speech separation benchmarks and robustness under reverberant conditions. Both single-stage TSE models and the two-stage separation model are trained with the negative SI-SDR loss using float16 mixed precision and a batch size of 24, as is the two-stage classification model.
All models were optimized using AdamW [29], with learning rates and gradient clipping configured individually for each model. The DPT-based single-stage TSE was trained with a learning rate of 1e-4, a maximum gradient norm of 30, and linear warm-up over 1,000 updates. The TF-Locoformer-based single-stage TSE was trained with a learning rate of 1e-4 and a maximum norm of 5, whereas the two-stage separation model used a learning rate of 1e-3 with the same maximum norm. The classification models were trained with a learning rate of 1e-4 and a maximum norm of 30. Different learning rates and maximum norm values were adopted to account for the distinct architectures and convergence behaviors of the respective models. The learning rate was halved if the validation loss did not improve for three consecutive epochs, and training employed early stopping with a maximum of 100 epochs.
IV-C Two-Stage Method versus Single-Stage Baselines
Experimental Setup: We evaluate the proposed two-stage method against single-stage baselines. In the first stage, the separation model is trained from scratch to obtain separated speech signals. In the subsequent classification stage, a single model is trained and evaluated using template prompts for each mixture, where each prompt corresponds to one of the fourteen relative cue types listed in Table III. The single-stage baselines include a DPT-based TSE model from our prior work [7] and a TF-Locoformer-based TSE model. The latter is derived by adapting the TF-Locoformer speech separation model for target speech extraction. Specifically, rather than estimating two masks (one per speaker) to separate both sources, the model is reconfigured to predict a single time–frequency mask corresponding exclusively to the target speaker. Samples labeled as “Similar” and “Same” categories are excluded from the training, validation, and test sets due to their lack of discriminative power. Additionally, we train and evaluate a classification model that uses an enrollment utterance from the target speaker to enable comparison with textual relative cues.
Results and Analysis:
TABLE V: Comparison of single-stage and two-stage TSE systems across different cues
Cue Type Single-Stage Two-Stage
DPT-based TSE TF-Locoformer TSE TF-Locoformer TSE
SI-SDRi PESQ SI-SDRi PESQ Sep ACC SI-SDRi PESQ
Random Cue $8.4$ $1.78$ $13.2$ $2.70$ 99.299.2% $17.1$ $3.39$
All Cue $8.9$ $1.80$ $14.3$ $2.88$ 99.899.8% $17.4$ $3.41$
Language $5.5$ $1.71$ $11.4$ $2.67$ 99.399.3% $17.3$ $3.41$
Transcription $4.3$ $1.67$ $8.7$ $2.47$ 90.490.4% $12.5$ $3.22$
Gender $8.8$ $1.80$ $12.8$ $2.75$ 98.198.1% $16.7$ $3.37$
Emotion $4.9$ $1.56$ $7.8$ $2.44$ 87.287.2% $10.3$ $3.10$
Pitch Level $7.9$ $1.77$ $12.8$ $2.75$ 94.694.6% $14.8$ $3.30$
Pitch Range $5.3$ $1.68$ $9.0$ $2.47$ 88.788.7% $11.3$ $3.14$
Loudness $8.7$ $1.80$ $13.7$ $2.84$ 97.297.2% $16.0$ $3.35$
Distance $8.3$ $1.78$ $14.6$ $2.89$ 98.498.4% $17.0$ $3.38$
Age Category 1.0-1.0 $1.38$ 1.9-1.9 $1.71$ 64.364.3% 1.6-1.6 $2.45$
Temporal Order $8.7$ $1.85$ $15.1$ $2.99$ 99.899.8% $18.0$ $3.45$
Speaking Rate $4.0$ $1.68$ $7.9$ $2.44$ 91.391.3% $13.1$ $3.27$
Speaking Duration $7.4$ $1.79$ $13.0$ $2.80$ 95.995.9% $15.6$ $3.34$
Enrollment Audio 95.795.7% $15.2$ $3.31$
We use scale-invariant signal-to-distortion ratio improvement (SI-SDRi) [24] and perceptual evaluation of speech quality (PESQ) [42] to evaluate target speech extraction performance. For the TF-Locoformer-based two-stage TSE framework, we report three metrics: the separation performance of TF-Locoformer (Sep SI-SDRi/PESQ), the classification accuracy on the separated signals (Sep ACC), and the overall target speech extraction performance (SI-SDRi/PESQ). The overall performance is calculated as the mean score between the predicted and ground-truth target speech under various relative cue conditions. The results are summarized in Table V.
To evaluate performance for different cue types, we follow a prompt-based setup: Individual cues: Each template prompt contains only a single relative cue type. For each cue type, the model is evaluated on mixtures corresponding to that cue, producing cue-specific SI-SDRi, PESQ, and classification metrics. All cue: Each prompt includes all available single relative cues presented in the target speech of a mixture. The model is evaluated on mixtures with multiple cue types simultaneously, allowing us to assess performance under the combination of all available cues. Random cue: Each prompt contains a random subset of relative cues, and evaluation is performed on mixtures corresponding to these randomly selected cues. This measures the robustness of the two-stage method under partial or unpredictable cue guidance.
The two-stage TSE also supports target speaker enrollment utterances as guidance, where the separated signal with the highest cosine similarity to the enrollment embedding is selected as the target. This enables a direct comparison with text-based relative cues. From the results, the TF-Locoformer-based single-stage TSE consistently outperforms the DPT-based single-stage TSE, achieving significant 3+ dB improvements in SI-SDRi and 0.7–1.1 improvements in PESQ, except for the age category, which remains challenging for both models. This suggests that age prediction performance is bottlenecked by the cue discriminability rather than the model architecture. The classification performance of different cue types reflects their relative contribution to the single-stage TSE baselines. Compared with the single-stage TF-Locoformer-based TSE, the two-stage method demonstrates clear advantages. It achieves over 3 dB SI-SDRi improvement for the random cue, all cue, language, transcription, gender, and speaking rate cues, and gains more than 2 dB SI-SDRi for the remaining cue types (except for the age category cue). Additionally, it provides an average improvement of over 0.5 in PESQ across all cue types, highlighting its strong practical potential.
Figure 3: Training and Validation Loss Comparison for TF-Locoformer-based single-stage TSE and TF-Locoformer-based Separation models over epochs. The loss is measured such that lower (more negative) is better.
Interestingly, some relative cue types—including all cue, random cue, language, gender, loudness, distance, temporal order and speaking duration—can even outperform the enrollment speech of the target speaker. This suggests that enrollment speech may contain potentially confounding information, as it could be recorded in a different environment and exhibit speech characteristics that differ from the target speech. These findings further highlight the practicality of using textual descriptions of relative cues, which not only address issues such as missing enrollment speech and privacy concerns but can also improve extraction performance by leveraging some distinguishable cue types present in the mixture speech samples.
We also present the training and validation loss curves for the TF-Locoformer-based single-stage TSE and TF-Locoformer-based separation models in Fig. 3. From the curves, we observe that the TF-Locoformer-based TSE model tends to overfit easily, whereas the TF-Locoformer-based separation model converges more gradually and achieves a better validation loss. In the early training epochs, the separation model also attains a better training loss, despite using a lower learning rate. These observations highlight the advantage of the two-stage approach.
To establish an oracle benchmark, we additionally train and evaluate a separate classification model on the target–inference reverberant–clean speech pairs from each mixture described in Section IV (Audio Mixtures of Dataset). The evaluation results are shown in Table VI.
TABLE VI: Performance across different cues
Cue Type Clean ACC Sep ACC
Random Cue 99.3% 99.2%
All Cue 99.8% 99.8%
Language 99.5% 99.3%
Transcription 90.8% 90.4%
Gender 98.0% 98.1%
Emotion 85.8% 87.2%
Pitch Level 95.2% 94.6%
Pitch Range 88.4% 88.7%
Loudness 98.8% 97.2%
Distance 99.0% 98.4%
Age Category 64.0% 64.3%
Temporal Order 100% 99.8%
Speaking Rate 91.5% 91.3%
Speaking Duration 96.3% 95.9%
Enrollment Audio 96.1% 95.9%
In general, classification performance on clean speech (Clean ACC) generally exceeds that on separated speech (Sep ACC), as expected. Interestingly, the accuracy for each cue type closely matches that of the model trained on separated speech generated by the TF-Locoformer separation model, demonstrating the high quality of the separated signals. The varying classification accuracies across cue types also provide insight into why some cues exert a stronger influence on target speech extraction, while others have a weaker effect.
Figure 4: Overall trend of classification accuracy with respect to the attribute differences between target and interfering speakers for each cue type.
We further analyze classification accuracy as a function of the continuous-valued attribute difference between paired target–interference clean speech samples for each cue type. The temporal order cue is excluded due to its perfect classification accuracy (100%). For the remaining cues, logistic regression is used to obtain a smooth estimate of overall classification accuracy over the continuous attribute space. As shown in Fig. 4, accuracy exhibits a clear increasing trend as attribute differences grow. Except for age category, all cues reach near-perfect classification when the differences become sufficiently large.
IV-D Relative Cues versus Independent Cues
Experimental Setup: We evaluate whether relative cues outperform independent cues for target speech extraction within a two-stage framework. For discrete-valued attributes, relative and independent cues provide the same level of discriminability; therefore, we restrict our evaluation to originally continuous-valued attributes. In total, eight cues are considered, of which two—age category and temporal order—are excluded. The age category cue is excluded because the training data are limited to speakers aged 18–50, making absolute age categorization unreliable. Applying common-sense age groupings (young: <25<25, youth: 25–40, middle: 41–64, senior: >65>65) would lead to severe class imbalance (youth: 10,310; young: 7,063; middle: 2,627; senior: 0). The temporal order cue, defined as the difference in onset time between the target and interfering speech, is inherently relational and therefore does not admit a meaningful absolute categorization. Consequently, the evaluation focuses on the remaining six cues: pitch level, pitch range, loudness, distance, speaking rate, and speaking duration.
Separated speech samples are selected from the TF-Locoformer separation model corresponding to these cues. Two binary classification models are trained: one guided by relative-cue text prompts and the other by independent-cue text prompts.
TABLE VII: Metrics for Independent and Relative Cues
Independent Cues Relative Cues
Cue Type Group N ACC SI-SDRi PESQ N ACC SI-SDRi PESQ
Overall non-same / non-similar 34501 93.5% $14.3$ $3.29$ 38974 94.1% $14.5$ $3.29$
Overall all 59009 75.3% $4.3$ $2.87$ 59009 79.2% $6.3$ $2.93$
Distance non-same / non-similar 5230 94.0% $14.4$ $3.29$ 2876 98.5% $17.0$ $3.38$
Distance all 10000 72.9% $2.9$ $2.81$ 10000 64.1% 2.1-2.1 $2.53$
Loudness non-same / non-similar 6994 91.8% $13.0$ $3.23$ 5008 97.9% $16.3$ $3.36$
Loudness all 10000 79.2% $6.1$ $2.94$ 10000 74.4% $3.7$ $2.85$
Pitch Level non-same / non-similar 6630 98.1% $16.8$ $3.39$ 8909 95.4% $15.2$ $3.32$
Pitch Level all 9990 82.4% $8.2$ $3.03$ 9990 90.4% $12.4$ $3.18$
Pitch Range non-same / non-similar 6652 89.7% $12.0$ $3.19$ 7612 89.6% $11.9$ $3.17$
Pitch Range all 9988 77.0% $5.0$ $2.90$ 9988 80.0% $6.5$ $2.93$
Speaking Rate non-same / non-similar 5979 90.8% $13.0$ $3.26$ 6796 89.8% $12.3$ $3.23$
Speaking Rate all 9031 75.9% $4.9$ $2.91$ 9031 80.1% $7.0$ $2.99$
Speaking Duration non-same / non-similar 3016 99.5% $18.6$ $3.50$ 7773 96.7% $16.0$ $3.36$
Speaking Duration all 10000 64.2% 1.2-1.2 $2.61$ 10000 86.3% $10.4$ $3.10$
Independent Cues
Independent cue categories are defined based on the distribution of attribute values in the training samples (including both target and interference speech), following previous studies[17, 18]. We use equal-frequency binning, where each attribute is quantized into two or three categories with similar amount of samples per each category:
  • Pitch level: low / normal / high
  • Pitch range: narrow / normal / wide
  • Loudness: quiet / normal / loud
  • Distance: near / far
  • Speaking rate: slow / normal / fast
  • Speaking duration: short / long
Relative Cues
Relative cue categories are defined by comparing the target speech with the interference speech:
  • Pitch level: lower / similar / higher
  • Pitch range: narrower / similar / wider
  • Loudness: quieter / similar / louder
  • Distance: nearer / similar / farther
  • Speaking rate: slower / similar / faster
  • Speaking duration: shorter / similar / longer
We use the audio data corresponding to six relative cues—pitch level, pitch range, loudness, distance, speaking rate, and speaking duration—from the dataset described in Section IV. To isolate the effect of cue category differences, we generate prompts for independent cues and regenerate prompts for relative cues using predefined templates. Each prompt describes a single cue, and all prompts share the same sentence structure, differing only in whether the cue is expressed in relative or independent terms. For example, a relative-cue prompt is “Please extract the speaker with a higher pitch level in the audio,” whereas the corresponding independent-cue prompt is “Please extract the speaker with a high pitch level in the audio.”
During training, samples in which both speech signals belong to the same absolute category (or the Similar relative category) are excluded. During evaluation, all samples are retained, and target labels are derived from the PIT results of the separation model.
Classification accuracy is reported under different conditions. For independent cues, three conditions are considered: (i) same category, (ii) adjacent categories (e.g., slow–normal for speaking rate), and (iii) distinctly different categories (e.g., slow–fast for speaking rate). For relative cues, two conditions are considered: (i) same category (“Similar”) and (ii) different categories. We further analyze cases in which target–interference pairs fall into different relative-cue categories but are quantized into the same, adjacent, or distinct absolute categories, highlighting scenarios where independent cues provide limited discriminability while relative cues remain informative.
(a) Independent Speaking Rate Cue
(b) Relative Speaking Rate Cue
Figure 5: Classification accuracy and corresponding sample distribution across regions for independent (a) and relative (b) speaking rate cues, where $x^{(\mathit{tar})}$ (y-axis) and $x^{(\mathit{inf})}$ (x-axis) represent the speaking rate values of the target and interfering speakers, respectively.
Results and Analysis: Overall classification accuracy is compared across cue types to evaluate the effectiveness of relative versus independent cues for TSE. To facilitate analysis, pairs are grouped as follows: for independent cues, pairs in the same category are labeled the same group, and pairs in different categories are labeled the non-same group; for relative cues, pairs in the same category are labeled the similar group, and pairs in different categories are labeled the non-similar group. Results are presented across six continuous-valued attributes.
For independent cues, we report the classification accuracy, as well as the resulting SI-SDRi and PESQ performance for the non-same and same groups. For relative cues, we report the classification accuracy and the corresponding SI-SDRi and PESQ performance for the non-similar and similar groups. We explicitly account for the number of samples (N) in each group to fairly compare whether relative cues can exploit more information than independent cues. Notably, each relative cue and its corresponding independent cue are evaluated using the same number of samples. The results are summarized in Table VII.
The overall classification accuracy, SI-SDRi, and PESQ achieved with relative cues surpass those obtained with independent cues. Moreover, the number of samples in the non-similar group is larger than that in the similar group, highlighting the advantage of relative cues in leveraging more discriminative information. In particular, for pitch level, pitch range, speaking rate, and speaking duration, the relative categories demonstrate a clear advantage over absolute categories in utilizing more distinguishing information.
For the distance and loudness cues, using relative categories results in more samples being assigned to the similar group than when using absolute categories. This is due to the limited ranges of the simulated speaker-to-microphone distances and SIR values. In our dataset, the speaker-to-microphone distances in the simulated mixtures range from 0.3 to 1.5 m, and an auditory distance threshold of 0.5 m is used to distinguish between the two speakers. Consequently, many target-interference speech pairs are classified as having similar distances. A similar effect is observed for the loudness cue: loudness is approximated using SIR, which is constrained to the range of -6 to 6 dB in our simulation. With a loudness cue threshold of 3 dB, nearly half of the samples are categorized as having similar loudness.
Figure 6: Comparison between independent and relative cue categorizations for the speaking rate attribute.
(a) Independent Speaking Rate Cue
(b) Relative Speaking Rate Cue
Figure 7: Joint distribution of classification accuracy for independent (a) and relative (b) speaking-rate cues across all paired target–interference samples. The x-axis and y-axis correspond to the speaking rates (syllables per minute) of the interfering and target speakers, respectively. Red points indicate correctly classified samples, while blue points indicate incorrectly classified samples. The gray region denotes target–interference speaking-rate differences within $\pm 15\%$.
It is informative to examine the specific categories of target–interference samples. Fig. 5 presents the classification accuracy and sample distribution for both relative and independent cues, using speaking rate as an example. When the target and interference speech fall into clearly distinct categories (e.g., slow–fast or fast–slow), the model achieves the highest accuracy. In contrast, accuracy decreases when they belong to adjacent categories (e.g., slow–normal or normal–fast), supporting the assumption in Section II.
To further analyze the relationship between independent and relative cue categorizations, we focus on target–interference pairs labeled as relative-different and examine whether they fall into the same, adjacent, or distinct absolute categories. Using speaking rate as an example (Fig. 6), we report classification accuracy under these three absolute conditions for both the full set and the relative-different subset, along with their sample distribution. Accuracy increases monotonically with the magnitude of speaking rate difference between two speakers. In the same condition, performance is relatively low (48.9%), but restricting to relative-different pairs improves accuracy to 79.4%, indicating that relative comparison provides additional discriminative information even when absolute categories are identical. In the adjacent condition, accuracy is high (87.3%) and slightly improves to 89.5% under the relative-different subset. In the distinct condition, performance reaches 98.3% in both cases. These results show that classification performance strongly correlates with speaking rate difference between two speakers, and that relative cues particularly benefit challenging cases where absolute quantization obscures pairwise distinctions.
Fig. 7 further illustrates the joint distribution of target and interference speaking rates under absolute and relative categorizations. Compared with independent cues, relative cues yield more correct classifications in the slow–slow, normal–normal, and fast–fast regions, highlighting their advantage when absolute categories are identical.
V Conclusion
We have presented a comprehensive study on text-guided target speech extraction that leverages inter-speaker relative cues across a wide range of speech attributes. By providing a theoretical analysis, redesigning the model architecture with a two-stage separation-and-classification approach, and empirically evaluating relative cues versus independent cues on comparable continue-valued attributes.
Experimental results show that the proposed two-stage approach substantially improves both signal-level (SI-SDRi) and perceptual (PESQ) performance, outperforming single-stage text-based TSE. Analysis of classification accuracy across cue types indicates that while some cues have strong discriminative power, certain text cues are even more effective than conventional audio-based methods, highlighting their practical potential. Finally, our results confirm that relative cues generally outperform independent cues, underscoring their advantage for text-guided target speech extraction. Notably, adapting a state-of-the-art separation model into a single-stage TSE system can also significantly enhance target speech extraction, further validating the effectiveness of leveraging advanced separation models in TSE frameworks
Overall, our findings underscore the effectiveness of integrating perceptually relative cues with a two-stage model for robust and flexible text-guided TSE. The proposed framework not only mitigates speaker confusion inherent in single-stage systems but also provides valuable insights into the relative importance of different speech attributes for TSE.

### v2

@@
-arXiv:2603.01316v1[eess.AS] 01 Mar 2026
+Inter-Speaker Relative Cues for Two-Stage Text-Guided Target Speech Extraction
 Abstract
 This paper investigates the use of relative cues for text-based target speech extraction (TSE).
-We first provide a theoretical justification for relative cues from the perspectives of human perception and label quantization, showing that relative cues preserve fine-grained distinctions often lost in absolute categorical representations.
-Building on this analysis, we propose a two-stage TSE framework, in which a speech separation model generates candidate sources, followed by a text-guided classifier that selects the target speaker based on embedding similarity.
-Using this framework, we train two separate classification models to evaluate the advantages of relative cues over independent cues in terms of both classification accuracy and TSE performance.
-Experimental results demonstrate that (i) relative cues achieve higher overall classification accuracy and improved TSE performance compared with independent cues, (ii) the two-stage framework substantially outperforms single-stage text-conditioned extraction methods on both signal-level and objective perceptual metrics, and (iii) certain relative cues (language, gender, loudness, distance, temporal order, speaking duration, random cue and all cue) can surpass the performance of an audio-based TSE system.
+We first provide a theoretical justification for relative cues from the perspectives of human perception and label quantization, showing that relative cues preserve fine-grained distinctions that are often lost in absolute categorical representations for continuous-valued attributes.
+Building on this analysis, we propose a two-stage TSE framework in which a speech separation model first generates candidate sources, followed by a text-guided classifier that selects the target speaker based on embedding similarity.
+Within this framework, we train two separate classification models to evaluate the advantages of relative cues over independent cues in case of continuous-valued attributes, considering both classification accuracy and TSE performance.
+Experimental results demonstrate that (i) relative cues achieve higher overall classification accuracy and improved TSE performance compared with independent cues; (ii) the proposed two-stage framework substantially outperforms single-stage text-conditioned extraction methods on both signal-level and objective perceptual metrics; and (iii) several relative cues, including language, loudness, distance, temporal order, speaking duration, random cues, and all cues, can even surpass the performance of an enrollment-audio-based TSE system.
 Further analysis reveals notable differences in discriminative power across cue types, providing insights into the effectiveness of different relative cues for TSE.
@@
 Subsequent developments incorporated pretrained speaker encoders, such as D-Vectors[49], X-Vectors[46], and ECAPA-TDNN[8], to obtain robust speaker embeddings, decoupling speaker modeling from the extraction network.
-Moreover, multi-modal approaches integrating audio, video cues have gained attention[37, 26, 54, 58], especially for audiovisual datasets like AVSpeech[11] or VoxCeleb[35].
+Moreover, multi-modal approaches integrating audio and video cues have gained attention[35, 24, 48, 52], especially for audiovisual datasets like AVSpeech[11] or VoxCeleb[34].
 These methods align visual speech movements or lip information with audio stream[3, 23], improving robustness in visually grounded TSE scenarios.
@@
 Beyond acoustic/visual-based guidance, recent research has explored text-guided extraction.
-These models use a text query (prompt) to describe speech attributes of target speaker as a condition to guide the extraction process of the desired speaker’s speech signal in the mixture, enabling extraction when no enrollment utterance of target speaker is available.
+These models use a text query (prompt) describing the speech attributes of a target speaker as a condition to guide the extraction of the desired speaker’s speech signal from a mixture, enabling extraction when no enrollment utterance of the target speaker is available.
 Compared to other cues, text offers greater accessibility, enabling users to provide instructions without recording audio or video.
 It also provides greater flexibility, as it can convey a wide range of target speaker characteristics, including linguistic cues (e.g., speech transcription), paralinguistic cues (e.g., gender), and acoustic cues (e.g., loudness of speech).
-In addition, text offers stronger privacy by avoiding the transmission or storage of personal voice data.
+In addition, text offers stronger privacy by avoiding the transmission or storage of voice data or other media.
 Hao et al.
@@
 Subsequently, Huo et al.
-proposed a flexible text-guided TSE framework and introduced the TextrolMix dataset, which provides rich textual descriptions of speech attributes, including speaker identity, emotion, pitch level, gender, accent, and speaking rate [17].
+proposed a flexible text-guided TSE framework and introduced the TextrolMix dataset, which provides rich textual descriptions of speech attributes, including speaker identity, emotion, pitch, gender, accent, and speaking rate [16].
 Similarly, Seki et al.
@@
 By formulating attributes relatively rather than as speaker-specific categories, these cues preserve fine-grained distinctions that are critical for target speaker extraction.
-To study these cues, we constructed a two-speaker mixture dataset spanning a wide range of attributes, including pitch level, pitch range, speaking rate, speaking duration, temporal order, age, loudness, speaker-to-microphone distance, language, gender, transcription and emotion [7].
+To study these cues, we constructed a two-speaker mixture dataset spanning a wide range of attributes, including pitch, pitch range, speaking rate, speaking duration, temporal order, age, loudness, speaker-to-microphone distance, language, gender, transcription and emotion [7].
 This approach is motivated by human perception: people perceive continuous-valued speech attributes primarily through comparative listening, rather than by measuring precise physical quantities.
-Instead of relying on exact values—such as fundamental frequency ($F_{0}$) in Hertz or speaking rate in syllables per minute—listeners judge whether one speaker sounds higher or lower in pitch, or faster or slower in speaking rate than another [53, 33, 32].
+Instead of relying on exact values, such as fundamental frequency ($F_{0}$) in Hertz or speaking rate in syllables per minute, listeners judge whether one speaker sounds higher or lower in pitch, or faster or slower in speaking rate than another [31, 30].
 Conventional TSE systems that rely on absolute, discretized labels fail to capture these relative distinctions.
 Subtle but perceptually meaningful differences between speakers can be lost when multiple speakers fall into the same category (e.g., both labeled as “fast” speaking rate).
-By explicitly encoding inter-speaker comparisons—such as “Speaker A has a faster speaking rate than Speaker B”—relative attribute cues preserve these contrasts, providing more discriminative and perceptually aligned conditioning for target speaker extraction.
+By explicitly encoding inter-speaker comparisons, such as “Speaker A has a faster speaking rate than Speaker B”, relative attribute cues preserve these contrasts, resulting in more discriminative and perceptually aligned conditioning for target speaker extraction.
+Although this study restricts mixtures to two speakers, the relative cue formulation can be naturally extended to mixtures with more speakers.
+For instance, in mixtures with three or more speakers, cues such as “fastest speaking rate” or “second-fastest speaking rate” can be used to specify the target speaker.
+For discrete-valued speech attributes, listeners identify the target speaker by comparing the attribute category with that of the interfering speaker, further illustrating the principle of relative cues rather than speaker-independent cues.
 Building upon these insights, the main contributions of this work are summarized as follows:
-  • Theoretical Justification of Relative Cues: We provide a theoretical analysis of relative cues, which accounts for human perception and label quantization, offering a principled explanation of their advantages over speaker independent cues, which we hereafter refer to as independent cues throughout the remainder of this paper.
-  • Proposed Two-Stage Text-Guided TSE Method: In our previous text-based TSE system—similar to conventional audio-based approaches—the conditioning-and-extraction paradigm is prone to speaker confusion [62, 28, 34, 10] during inference, leading to degraded performance.
-Motivated by enrollment-audio-based two-stage TSE approach [10], and to improve upon our previous single-stage system, we adopt a two-stage framework in this work: first, performing speech separation, and then using a classification model to select the target speaker based on inter-speaker relative cues embedded in text prompts.
+  • Theoretical Justification of Relative Cues: We provide a theoretical analysis of relative cues, which accounts for human perception and label quantization, offering a principled explanation of their advantages over speaker-independent cues, which we hereafter refer to as independent cues throughout the remainder of this paper.
+  • Proposed Two-Stage Text-Guided TSE Method: Similar to conventional audio-based approaches, our previous text-based TSE system follows a conditioning-and-extraction paradigm that is prone to speaker confusion during inference [55, 26, 33, 10], leading to degraded performance.
+Motivated by enrollment-audio-based two-stage TSE approach [10], and to improve upon our previous single-stage system, we adopt a two-stage framework in this work: first, performing speaker-independent speech separation, and then using a classification model to select the target speaker based on inter-speaker relative cues embedded in text prompts.
 This approach not only mitigates the speaker confusion inherent in single-stage TSE systems but also provides interpretability, allowing insight into the relative effectiveness of different cues through analysis of the classification accuracy for each cue.
-  • Empirical Evaluation of the Advantages of Relative Cues: Based on the proposed two-stage framework, we evaluate the impact of using relative versus independent cues on overall target speech classification and the corresponding TSE performance, focusing on continuous-valued attributes.
-This provides an empirical perspective on the advantages of relative cues over independent cues.
+  • Empirical Evaluation of the Advantages of Relative Cues: Based on the proposed two-stage framework, we evaluate the impact of relative versus independent cues on overall target speaker classification and the corresponding TSE performance for comparable continuous-valued attributes.
+This analysis provides an empirical perspective on the advantages of relative cues over independent cues.
 The remainder of this paper is organized as follows.
@@
 II Inter-Speaker Relative Cues
+(a) Independent Cue
+(b) Relative Cue for Direct Difference
+(c) Relative Cue for Percentage Difference
+Figure 1: Illustration of quantization schemes for continuous-valued attributes.
+(a) Independent cues: the attribute value space of the target and interfering speakers, denoted by $x_{i}^{(\mathit{tar})}$ (y-axis) and $x_{i}^{(\mathit{inf})}$ (x-axis), is partitioned into a small number of coarse categories (three shown as in the example).
+Shaded regions indicate attribute values of the two speakers that are quantized into the same category, in which case the two speakers cannot be distinguished based on the attribute.
+(b) Relative cues for direct differences: attribute values of the target and interfering speakers are compared with direct difference.
+The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold.
+Regions outside the shaded area represent perceptually distinct differences.
+(c) Relative cues for percentage differences: attribute values of the target and interfering speakers are compared with percentage difference.
+The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold.
+Because the comparison is based on relative differences, the Similar region gradually expands from the zero point.
+Regions outside the shaded area represent perceptually distinct relative differences.
 The proposed text-guided TSE system takes a two-speaker mixture as input along with a text prompt which consists of cues indicating the target speaker, aiming to extract the target speech signal.
@@
   8. 8.
-    appearance time, defined as the onset time of a speaker’s speech within the mixture.
+    appearance time, defined as the onset time of the first occurrence of a speaker’s speech in the mixture.
 Discrete-valued speech attributes correspond to categorical properties that take values from a finite set. In this study, these attributes include:
@@
   4. 4.
-    transcription content, corresponding to the linguistic content of the utterance in textual form.
-We derive each inter-speaker relative cues by comparing a single speech attribute between the target speaker and the interfering speaker in a two-speaker mixture.
+    transcription, corresponding to the linguistic content of the utterance in textual form.
+We derive each inter-speaker relative cue by comparing a single speech attribute between the target speaker and the interfering speaker in a two-speaker mixture.
 Each relative cue encodes the difference along one specific attribute only.
@@
 $\tilde{x}_{i}=\mathcal{Q}_{i}(x_{i}),\quad\tilde{x}_{i}\in\{\mathit{A}_{i,1},\mathit{A}_{i,2},\dots,\mathit{A}_{i,k_{i}}\},$ (1)
-where $\mathcal{Q}_{i}(\cdot)$ denotes the quantization operator for the $i$-th attribute, $k_{i}$ is the number of discrete categories for that attribute, and $\mathit{A}_{i,1},\dots,\mathit{A}_{i,k_{i}}$ represent the discrete categories (e.g., low, normal, and high for mean $F_{0}$).
-The corresponding independent cue category for attribute $i$ is given by $\tilde{x}_{i}$.
-This mapping from continuous-valued attribute values $x_{i}$ to discrete independent cue categories $\tilde{x}_{i}$ is illustrated in Fig.
+where $\mathcal{Q}_{i}(\cdot)$ denotes the quantization operator for the $i$-th attribute, $k_{i}$ is the total number of discrete categories for that attribute, and $\mathit{A}_{i,1},\dots,\mathit{A}_{i,k_{i}}$ denote the corresponding categories (e.g., low, normal, and high for mean $F_{0}$).
+The independent cue category for attribute $i$ is given by $\tilde{x}_{i}$.
+The mapping from continuous-valued attribute values $x_{i}$ to discrete independent cue categories $\tilde{x}_{i}$ is illustrated in Fig.
 1(a).
@@
 1(a)), in which case the attributes of two speakers cannot be distinguished from each other.
-(a) Independent Cue
-(b) Relative Cue for Absolute Difference
-(c) Relative Cue for Percentage Difference
-Figure 1: Illustration of quantization schemes for continuous-valued attributes.
-(a) Independent cues: the attribute value space of the target and interfering speakers, denoted by $x_{i}^{(\mathit{tar})}$ (y-axis) and $x_{i}^{(\mathit{inf})}$ (x-axis), is partitioned into a small number of coarse categories (three shown as in the example).
-Shaded regions indicate attribute values of the two speakers that are quantized into the same category, in which case the two speakers cannot be distinguished based on the attribute.
-(b) Relative cues for absolute differences: attribute values of the target and interfering speakers are compared with absolute difference.
-The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold.
-Regions outside the shaded area represent perceptually distinct absolute differences.
-(c) Relative cues for percentage differences: attribute values of the target and interfering speakers are compared with percentage difference.
-The shaded Similar region corresponds to $|\Delta x_{i}^{(*)}|\leq\theta_{i}$, indicating differences below the discrimination threshold.
-Because the comparison is based on relative differences, the Similar region gradually expands from the zero point.
-Regions outside the shaded area represent perceptually distinct relative differences.
 II-B Relative Cues
 In contrast, relative cues categorize continuous-valued attributes into groups (e.g., slower, similar, and faster for speaking rate) by comparing the attribute values of the two speakers.
-For the attributes RMS energy (dB), speaker-to-microphone distance, speaker age, and appearance time in the mixture, the difference is expressed as the absolute difference
-$\Delta x_{i}^{|.|}=x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}.$ (2)
-These attributes correspond to interpretable physical or temporal quantities whose perceptual relevance is largely determined by their absolute values.
-Changes in these quantities can be directly perceived and interpreted by human listeners without reference to a baseline.
+For the attributes RMS energy (dB), speaker-to-microphone distance, speaker age, and appearance time in the mixture, the difference is expressed as the direct difference
+$\Delta x_{i}^{(.)}=x_{i}^{(\mathrm{tar})}-x_{i}^{(\mathrm{inf})}.$ (2)
+These attributes correspond to interpretable physical or temporal quantities whose perceptual relevance is largely determined by differences in their values.
+In human perception, changes in stimulus magnitude can be detected through difference thresholds, allowing listeners to perceive relative differences in such attributes between competing sources.
+Consequently, variations in these quantities can be directly perceived and interpreted without requiring an explicit baseline reference [32, 6].
 For attributes such as mean $F_{0}$, $F_{0}$ span, speaking rate, and speaking duration, we employ the percentage difference
@@
 These attributes represent global prosodic properties, for which human listeners are more sensitive to proportional than absolute changes [12, 21, 47], making this formulation better aligned with human auditory perception across different baselines.
-For notational convenience, we introduce the variable $\Delta x_{i}^{(*)}\in\{\Delta x_{i}^{|.|},\Delta x_{i}^{\%}\}$ to denote the attribute-specific difference measure.
+For notational convenience, we introduce the variable $\Delta x_{i}^{(*)}\in\{\Delta x_{i}^{(.)},\Delta x_{i}^{\%}\}$ to denote the attribute-specific difference measure.
 This measure is then compared against the predefined threshold $\theta_{i}$ of each attribute, which are shown in Table I.
-TABLE I: Continuous-Valued Speech Attributes with Our Defined Thresholds for Relative Cues and Motivated Speech Perceptual Studies
-Speech Attribute Threshold ($\theta_{i}$) Motivated Findings
+TABLE I: Continuous-Valued Speech Attributes with Our Defined Thresholds for Relative Cues and Motivating Speech Perceptual Studies
+Speech Attribute Threshold ($\theta_{i}$) Findings from Motivating Speech Perception Literature
 RMS energy (dB) 3 dB Perception of relative loudness differences in speech is mediated by changes in loudness-related acoustic features, such as excitation strength derived from the glottal source [43].
@@
 [28] report that the average just-noticeable difference (JND) in speech-to-noise ratio is approximately 3 dB.
-Signal-to-interference ratio (SIR), computed based on RMS energy over active speech regions, can serve as an approximation of the perceived loudness difference between target and interfering speech.
 Speaker-to-microphone distance 0.5 m Human perception of speaker-to-microphone distance relies on relative variations in distance-dependent acoustic cues, particularly changes in the direct-to-reverberant energy ratio [44, 29, 53].
@@
 1(c).
-The distribution of these relative cues under different threshold settings in the test set (based on the simulated dataset described in Section IV) are presented in Table II.
-The total number of samples is almost balanced across cue types, except for speaker age, for which labeled samples are limited in our dataset.
-TABLE II: Sample counts of relative categories for each continuous attribute in the test set, based on the defined thresholds.
-Continue-valued Attribute R1 Similar R2
+The distribution of these relative cues under different threshold settings in the test set (based on the simulated dataset described in Section IV) is presented in Table II.
+The total number of samples is nearly balanced across cue types, except for speaker age, for which labeled samples are limited in our dataset.
+TABLE II: Sample counts of relative categories for each continuous attribute in the test set based on defined thresholds
+Continuous-valued Attribute R1 Similar R2
 RMS energy 2547 4992 2461
@@
 II-D Text Prompt Generation
-Inter-speaker relative cues are formulated as natural-language descriptions (text prompts) that explicitly characterize the relative differences between the target and interfering speakers.
-These prompts are used for both training and evaluation of the TSE systems.
-Each text prompt is generated by populating predefined templates with one or more relative cue descriptors, following the same approach as [7].
-The templates incorporate action-oriented verbs such as extract, separate, or isolate, combined with descriptive phrases that emphasize the distinguishing characteristics of the target speaker.
-For example, a cue vector indicating (female, higher pitch, faster speaking rate) can be verbalized as: “Please extract the female speaker with a higher pitch and faster speaking rate.” We also employed the GPT-4o-mini model to automatically rephrase the templates into more generic prompts.
-Although these diversified prompts increased linguistic variability, they resulted in lower performance.
-Therefore, we adopt the original template-based prompts in our experiments.
-Since relative cues are mixture-specific, the target speaker in a given mixture may be specified by one or multiple relative cues.
-Table IV reports the number of mixtures in each subset that contain a given relative cue.
-Since a single mixture may include multiple cues, the counts for individual cues do not sum to the total number of mixtures.
-As shown in the table, the target speaker in each mixture can be specified using different types of prompts.
-The Random Cue type consists of a combination of randomly selected relative cues, where the number of selected cues ranges from 2 to (n1)(n-1) and $n$ denotes the total number of relative cues in the complete set.
-In contrast, the All Cue type includes all $n$ relative cues for a given mixture.
-This procedure yields a diverse set of natural-language descriptions that accurately reflect the underlying attribute differences between speakers.
+Inter-speaker relative cues are formulated as natural language descriptions (i.e., text prompts) that explicitly capture the differences between the target and interfering speakers, and are used for both training and evaluation of the TSE systems.
+For each mixture, a target speaker is specified based on the dataset annotations (see Section IV), and a corresponding prompt is generated by instantiating a predefined template with one or more relative cues.
+Specifically, one prompt is generated for each available individual cue, as well as for combinations of cues, either randomly selected or comprising all available cues (see details below).
+In the first experimental setting (Section IV-C), cue categories are chosen to ensure clear inter-speaker contrast, thereby providing sufficient discriminability for target speaker extraction.
+In contrast, the second experimental setting (Section IV-D) adopts a different definition of cue categories, allowing labels such as “same” or “similar” during testing.
+The template structure used for prompt generation follows the implementation described in the code [^2] https://github.com/daiwangsnr/Inter-Speaker-Relative-Cues-for-Text-Guided-Target-Speech-Extraction/blob/main/prompt_template.py, incorporating action-oriented verbs such as extract, separate, or isolate, along with descriptive phrases that emphasize the distinguishing characteristics of the target speaker.
+For example, a cue vector such as (female, higher pitch, faster speaking rate) can be verbalized as: “Please extract the female speaker with a higher pitch and faster speaking rate.”
+Prompts are generated in three configurations for each mixture:
+  • Individual cues: A separate prompt is generated for each available cue.
+  • Random cues: If at least three cues are available, a subset is randomly sampled for each mixture, with size $2$ to n1n-1 (where $n>3$ is the total number of available cues), and used to generate a prompt.
+This configuration enables robust evaluation under partial and randomly combined cues.
+  • All cues: If at least two cues are available, a prompt is generated that incorporates all cues, enabling evaluation under fully combined cue conditions.
+We also experimented with GPT-4o-mini to automatically rephrase templates for greater linguistic diversity.
+Although this increased variation, performance slightly decreased; thus, template-based prompts are retained in our experiments.
+Table III (Section IV) reports the number of mixtures associated with each relative cue in each subset.
+Some cue names differ from their corresponding speech attributes (e.g., loudness [^3] In this work, “loudness” refers to RMS energy and does not strictly correspond to its conventional perceptual definition vs.
+RMS energy, age vs.
+speaker age, pitch vs.
+mean $F_{0}$, pitch range vs.
+$F_{0}$ span, distance vs.
+speaker-to-microphone distance, temporal order vs.
+appearance time, and emotion vs.
+emotional state).
+These differences arise because some cue names are abbreviated, while others are expressed in more perceptually meaningful terms.
+Since not all mixtures include every cue (i.e., the target speaker is not necessarily characterized by all possible cues), the counts for individual cues do not sum to the total number of mixtures.
+Figure 2: Overview of the proposed two-stage text-guided TSE system during inference.
 III Proposed Two-Stage Method
-Figure 2: Overview of the proposed two-stage text-guided TSE system during inference.
 In this work, we adopt a two-stage approach for text-guided target speech extraction, comprising: (1) speech separation and (2) target speech classification guided by a text prompt.
@@
 Stage 1: Speech Separation
-In the first stage, the mixture is separated into individual speech signals using a speech separation model $\mathcal{S}(\cdot)$
-$\hat{s}_{1},\hat{s}_{2}]=\mathcal{S}(\text{mix}),$
+In the first stage, the mixture is separated into individual speech signals using a speech separation model $\mathcal{S}(\cdot)$ as
+$\hat{s}_{1},\hat{s}_{2}]=\mathcal{S}({mix}),$
 trained with utterance-level permutation invariant training (PIT) [20], where $\hat{s}_{1}$ and $\hat{s}_{2}$ denote the estimated speech signals of the individual speakers.
@@
 $\tilde{z}_{j}=\mathcal{A}(\hat{s}_{j}).$
-The audio encoder consists of a self-supervised speech model (wav2vec2-large-xlsr-53222https://huggingface.co/facebook/wav2vec2-large-xlsr-53), followed by a pooling operation, ReLU activation, and layer normalization.
+The audio encoder consists of a pre-trained self-supervised speech model (wav2vec2-large-xlsr-53444https://huggingface.co/facebook/wav2vec2-large-xlsr-53), followed by a pooling operation, ReLU activation, and layer normalization.
 Similarly, the text prompt $p$ is encoded into the same embedding space using a text encoder $\mathcal{T}$ as
@@
 $\text{sim}_{j}=\frac{\tilde{z}_{p}\cdot\tilde{z}_{j}}{\|\tilde{z}_{p}\|\,\|\tilde{z}_{j}\|},\quad j\in\{1,2\}.$
-We define the logit as the difference between these similarity scores as
+We define the logit as the difference between the two similarity scores as
 $\text{logit}=\text{sim}_{1}-\text{sim}_{2},$
-which represents how much more similar the first separated signal $\hat{s}_{1}$ is to the text prompt compared with the second signal $\hat{s}_{2}$.
-The logit is then mapped to a probability using the sigmoid function:
-$\hat{y}=\sigma(\text{logit}),$
-where $\hat{y}\in[0,1]$ indicates the likelihood that $\hat{s}_{1}$ corresponds to the target speaker described by the text prompt. Finally, the predicted target speaker index is determined as
+which measures how much more similar the separated signal $\hat{s}_{1}$ is to the text prompt compared to $\hat{s}_{2}$.
+The logit is scaled by a temperature parameter $T$ ($T<1$) and mapped to a probability using the sigmoid function:
+$\hat{y}=\sigma(\text{logit}/T),$
+where $\hat{y}\in[0,1]$ represents the likelihood that $\hat{s}_{1}$ corresponds to the target speaker described by the prompt.
+A temperature $T<1$ sharpens the output distribution, making predictions more confident (closer to 0 or 1).
+Empirically, we find that $T=0.2$ yields good performance.
+The predicted target speaker index is then given by
 $\text{pred}=\begin{cases}1,&\hat{y}>0.5,\\ 0,&\text{otherwise}.\end{cases}$
-During inference, the signal with the higher similarity is selected as the target speech $\hat{s}_{\text{tar}}$.
+During inference, the signal with the higher similarity score is selected as the target speech $\hat{s}_{\text{tar}}$.
 Training Label Generation and Loss
@@
 IV Evaluation
-The evaluation aims to: (i) assess the advantages of the proposed two-stage framework over conventional single-stage systems, and (ii) examine whether relative cues yield better performance than independent cues.
+The evaluation aims to: (i) assess the advantages of the proposed two-stage framework over conventional single-stage systems and investigate the relative effectiveness of different cues within the two-stage framework, and (ii) examine whether relative cues yield better performance than independent cues.
 The following sections provide an overview of the experimental dataset and training configurations, followed by detailed descriptions of each experiment, including the setup, results, and analysis.
 IV-A Dataset
-To ensure broad linguistic and acoustic diversity, audio mixtures were constructed from multiple publicly available corpora spanning five major languages—Chinese, English, French, German, and Spanish—each providing annotated attributes such as language, gender, emotion, age, and transcription (see Table II).
+To ensure broad linguistic and acoustic diversity, we construct audio mixtures using the same corpora as in our prior work[7].
 Each mixture contains speech from two speakers—one designated as the target and the other as the interference—together with a text prompt describing their relative characteristics.
@@
 These attributes serve as the basis for constructing inter-speaker relative cues (e.g., higher pitch, faster speaking rate).
-All attributes are computed from the isolated clean speech signals prior to mixing, ensuring that the measurements are not affected by interference from other speakers or by mixture-level acoustic artifacts.
+All attributes are computed from the isolated clean speech signals prior to mixing, ensuring that the measurements are not affected by interference from other speakers or reverberation.
 Mean $F_{0}$ and $F_{0}$ span are extracted using the Librosa pYIN library [^6] https://librosa.org/doc/main/generated/librosa.pyin.html.
@@
 Syllable counts are estimated based on vowel sequences: in Chinese, each character is treated as one syllable, whereas in English, French, German, and Spanish, syllables are approximated by counting contiguous vowel groups within words.
-The RMS energy is computed from the reverberant speech while accounting for the speaking duration.
-The speaker-to-microphone distance is randomly sampled in a virtual room, with horizontal distances of 0.3–1.5 m to the microphone.
-Language, gender, age, emotional state, and transcription content are obtained directly from the original dataset annotations.
-(2) Audio Mixtures: Each corpus is split into non-overlapping training, validation, and test subsets, with no overlap in speakers or spoken words.
+The RMS energy is computed from reverberant speech, averaged over active speech segments to account for speaking duration.
+Since the mixtures include room acoustic simulations, the speaker-to-microphone distance is randomly sampled in a virtual room, with horizontal distances of 0.3–1.5 m to the microphone.
+Language, gender, age, emotional state, and transcription are obtained directly from the original dataset annotations.
+(2) Audio Mixtures: Each corpus is divided into non-overlapping training, validation, and test subsets, with disjoint speakers and no shared utterances.
 We generated 100,000 mixtures for training and 10,000 mixtures each for validation and testing.
 To create realistic speech mixtures, each mixture is composed of two speech signals, $S_{1}$ and $S_{2}$, from different speakers.
-Both signals are first processed to remove silence—using forced alignment for transcribed speech or voice activity detection for untranscribed segments—and then temporally arranged with partial overlaps to mimic conversational dynamics.
+Both signals are first processed to remove initial and trailing silence, using forced alignment for transcribed speech or voice activity detection for untranscribed segments, and are subsequently temporally arranged with partial overlaps to mimic conversational dynamics.
 If either $S_{1}$ or $S_{2}$ is shorter than 3 s, the overlap duration is set to the length of the shorter signal, and zero-padding is applied at the beginning of the shorter signal using a randomly selected start offset uniformly drawn from $0,\text{duration of longer sample}-\text{duration of shorter sample}]$.
 For longer signals, the overlap duration is set to $(\text{length}(S_{1})+\text{length}(S_{2})-6)$ s, with $S_{2}$ padded using a fixed start offset of $6-\text{length}(S_{2})$ s.
-To incorporate realistic room acoustics, each $S_{1}$ and $S_{2}$ is convolved with a separate room impulse response (RIR) from the gpuRIR library [9], simulating different speaker positions within the room.
-The simulated rooms vary in size, reverberation time, and speaker–microphone distance.
-Room dimensions are randomly sampled within 9–11 m in length and width, and 2.6–3.5 m in height, while reverberation times (RT60) are sampled from 0.3–0.6 s.
-The microphone is positioned at the center of the room, and speaker locations are randomly assigned such that horizontal speaker–microphone distances range from 0.3 to 1.5 m, with speaker heights between 1.6 and 1.9 m.
-We keep the reverberant $S_{2}$ fixed and adjust the reverberant $S_{1}$ by applying a signal-to-interference ratio (SIR), uniformly sampled from –6 to 6 dB.
-For each mixture, one of the reverberant signals, $S_{1}$ or $S_{2}$, is randomly selected to serve as the target speech.
-TABLE III: Summary of used speech datasets across five languages, detailing the number of speakers in the training, validation, and test splits, along with their partial attributes.
-Dataset Language Speakers Attributes
-CASIA [60] Chinese 2/1/1 Language, Emotion, Gender, Transcription
-ESD [63] Chinese 6/2/2 Language, Emotion, Gender, Transcription
-Aishell3 [49] Chinese 120/30/30 Language, Gender, Transcription
-MagicData-Read [36] Chinese 33/8/11 Language, Age, Gender
-Librispeech [38] English 120/30/30 Language, Gender, Transcription
-ESD [63] English 6/2/2 Language, Emotion, Gender, Transcription
-OreauFR-02 [21] French 21/7/4 Language, Emotion, Age, Gender
-MLS French [39] French 90/18/18 Language, Gender, Transcription
-EmoDB [6] German 6/2/2 Language, Emotion, Age, Gender, Transcription
-MLS German [39] German 120/30/30 Language, Gender, Transcription
-MLS Spanish [39] Spanish 79/20/20 Language, Gender, Transcription
-EmoMatchSpanish [14] Spanish 30/10/10 Language, Emotion, Gender, Transcription
-(3) Text Prompts: As detailed in Section II, Table IV summarizes the distribution of samples across relative cue categories in the training, validation, and test sets, excluding the “Similar” and “Same” category, as they do not offer discriminative power.
-TABLE IV: Sample counts of each relative cue in the training, validation and test sets
+To simulate realistic room acoustics, each $S_{1}$ and $S_{2}$ is convolved with a separate room impulse response (RIR) from the gpuRIR library[9], representing different speaker positions.
+Room dimensions are randomly sampled within 9–11 m (length/width) and 2.6–3.5 m (height), with RT60 values of 0.3–0.6 s.
+The microphone is positioned at the room center, and speakers are placed such that horizontal speaker–microphone distances range from 0.3 to 1.5 m and heights from 1.6 to 1.9 m.
+The reverberant $S_{2}$ is fixed while $S_{1}$ is scaled to a signal-to-interference ratio (SIR) uniformly sampled between –6 and 6 dB.
+For each mixture, one of the reverberant signals is randomly selected as the target speech.
+(3) Text Prompts: The details of text prompts generation are described in Section II, Table III summarizes the distribution of samples across relative cue categories in the training, validation, and test sets, excluding the “Similar” and “Same” category, as they do not offer discriminative power.
+Since the “random” and “all” cues are not available in every mixture to specify the target speaker, the number of samples associated with these cues does not correspond to the total number of mixtures.
+TABLE III: Sample counts of each relative cue in the training, validation, and test sets, excluding cues labeled as “Similar” or “Same”
 Relative Cue Training Validation Test
-Random Cue 99641 9962 9956
-All Cue 99960 9996 9997
+Random Cues 99641 9962 9956
+All Cues 99960 9996 9997
 Language 73845 7192 7157
@@
 Speaking Duration 76498 7880 7773
-Pitch Level 89366 8917 8909
+Pitch 89366 8917 8909
 Pitch Range 77019 7715 7612
 Distance 27983 2672 2876
-Age Category 3130 348 395
+Age 3130 348 395
 Loudness 50731 4971 5008
@@
 IV-B Training Configurations
-For the speech encoder used in the classification model of the proposed two-stage framework, all cue types employed during training span linguistic, paralinguistic, and acoustic information.
-Therefore, the speech encoder plays a central role in capturing these diverse representations.
+For the speech encoder in the classification model of the proposed two-stage framework, all cue types span linguistic, paralinguistic, and acoustic information, making the encoder central to capturing these diverse representations.
 We adopt the encoder from the pretrained wav2vec2-large-xlsr-53 model, a multilingual self-supervised speech representation model based on the wav2vec 2.0 architecture [4].
-Trained on speech data from 53 languages, including Multilingual LibriSpeech and other large-scale corpora, wav2vec2-large-xlsr-53 learns universal, language-independent representations directly from raw audio.
-Compared to the state-of-the-art speaker encoder ECAPA-TDNN, which is primarily designed to extract speaker-specific acoustic embeddings, wav2vec2-large-xlsr-53 provides richer representations that jointly capture linguistic, paralinguistic, and acoustic information.
-This makes it particularly well suited for cross-lingual target speech extraction guided by diverse cue types.
-For wav2vec2-large-xlsr-53, only the first five encoder layers are used, with fine-tuning applied to the fifth layer.
-To encode text prompts into embeddings, we employ the LLaMA 3.2 1B Instruct model for all single-stage TSE systems as well as for the classification model in the two-stage framework.
-The LLaMA 3.2 1B instruction model was fine-tuned using LoRA (rank 16, scaling factor 16, dropout 0.05) on the query and key projection layers of self-attention.
-To ensure high-quality signal separation, we adopt the SOTA speech separation model TF-Locoforme [43], which has demonstrated superior performance on public speech separation benchmarks and robustness under reverberant conditions.
-Both single-stage TSE models and the two-stage separation model are trained with the negative SI-SDR loss using float16 mixed precision and a batch size of 24, as is the two-stage classification model.
-All models were optimized using AdamW [29], with learning rates and gradient clipping configured individually for each model.
-The DPT-based single-stage TSE was trained with a learning rate of 1e-4, a maximum gradient norm of 30, and linear warm-up over 1,000 updates.
+Trained on speech from 53 languages, including Multilingual LibriSpeech and other large-scale corpora, it learns universal, language-independent representations directly from raw audio.
+Compared to the ECAPA-TDNN speaker encoder, which focuses on speaker-specific embeddings, wav2vec2-large-xlsr-53 captures richer linguistic, paralinguistic, and acoustic features, making it well suited for cross-lingual target speech extraction.
+We use the first five encoder layers, with fine-tuning applied to the fifth layer.
+To encode text prompts into embeddings, we use the LLaMA 3.2 1B Instruct model for both single-stage TSE systems and the classification model in the two-stage framework.
+The LLaMA 3.2 1B instruction model is fine-tuned with LoRA (rank 16, scaling factor 16, dropout 0.05) on the query and key projection layers of self-attention.
+To ensure high-quality signal separation, we adopt the TF-Locoformer [39], which demonstrates state-of-the-art performance and robustness under reverberation.
+Single-stage TSE and two-stage separation models are trained with negative SI-SDR loss, float16 mixed precision, and a batch size of 24 (24 V100 GPUs, batch size 1 per GPU).
+The classification models also use float16 precision and batch size 24.
+All models are optimized using AdamW [27], with learning rates and gradient clipping configured individually for each model.
+The dual-path transformer (DPT)-based single-stage TSE model (used in the following subsection) was trained with a learning rate of 1e-4, a maximum gradient norm of 30, and 1000-step linear warm-up.
 The TF-Locoformer-based single-stage TSE was trained with a learning rate of 1e-4 and a maximum norm of 5, whereas the two-stage separation model used a learning rate of 1e-3 with the same maximum norm.
-The classification models were trained with a learning rate of 1e-4 and a maximum norm of 30.
+The classification model use a learning rate of $1\times 10^{-4}$ with max norm 30.
 Different learning rates and maximum norm values were adopted to account for the distinct architectures and convergence behaviors of the respective models.
-The learning rate was halved if the validation loss did not improve for three consecutive epochs, and training employed early stopping with a maximum of 100 epochs.
-IV-C Two-Stage Method versus Single-Stage Baselines
-Experimental Setup: We evaluate the proposed two-stage method against single-stage baselines.
-In the first stage, the separation model is trained from scratch to obtain separated speech signals.
-In the subsequent classification stage, a single model is trained and evaluated using template prompts for each mixture, where each prompt corresponds to one of the fourteen relative cue types listed in Table III.
-The single-stage baselines include a DPT-based TSE model from our prior work [7] and a TF-Locoformer-based TSE model.
-The latter is derived by adapting the TF-Locoformer speech separation model for target speech extraction.
-Specifically, rather than estimating two masks (one per speaker) to separate both sources, the model is reconfigured to predict a single time–frequency mask corresponding exclusively to the target speaker.
-Samples labeled as “Similar” and “Same” categories are excluded from the training, validation, and test sets due to their lack of discriminative power.
-Additionally, we train and evaluate a classification model that uses an enrollment utterance from the target speaker to enable comparison with textual relative cues.
-Results and Analysis:
-TABLE V: Comparison of single-stage and two-stage TSE systems across different cues
+Learning rates are halved if validation loss does not improve for three epochs, and training employs early stopping with a maximum of 100 epochs.
+TABLE IV: Comparison of single-stage and two-stage TSE systems across different cues
 Cue Type Single-Stage Two-Stage
@@
 Emotion $4.9$ $1.56$ $7.8$ $2.44$ 87.287.2% $10.3$ $3.10$
-Pitch Level $7.9$ $1.77$ $12.8$ $2.75$ 94.694.6% $14.8$ $3.30$
+Pitch $7.9$ $1.77$ $12.8$ $2.75$ 94.694.6% $14.8$ $3.30$
 Pitch Range $5.3$ $1.68$ $9.0$ $2.47$ 88.788.7% $11.3$ $3.14$
@@
 Distance $8.3$ $1.78$ $14.6$ $2.89$ 98.498.4% $17.0$ $3.38$
-Age Category 1.0-1.0 $1.38$ 1.9-1.9 $1.71$ 64.364.3% 1.6-1.6 $2.45$
+Age 1.0-1.0 $1.38$ 1.9-1.9 $1.71$ 64.364.3% 1.6-1.6 $2.45$
 Temporal Order $8.7$ $1.85$ $15.1$ $2.99$ 99.899.8% $18.0$ $3.45$
@@
 Speaking Duration $7.4$ $1.79$ $13.0$ $2.80$ 95.995.9% $15.6$ $3.34$
-Enrollment Audio 95.795.7% $15.2$ $3.31$
-We use scale-invariant signal-to-distortion ratio improvement (SI-SDRi) [24] and perceptual evaluation of speech quality (PESQ) [42] to evaluate target speech extraction performance.
-For the TF-Locoformer-based two-stage TSE framework, we report three metrics: the separation performance of TF-Locoformer (Sep SI-SDRi/PESQ), the classification accuracy on the separated signals (Sep ACC), and the overall target speech extraction performance (SI-SDRi/PESQ).
-The overall performance is calculated as the mean score between the predicted and ground-truth target speech under various relative cue conditions.
-The results are summarized in Table V.
-To evaluate performance for different cue types, we follow a prompt-based setup: Individual cues: Each template prompt contains only a single relative cue type.
-For each cue type, the model is evaluated on mixtures corresponding to that cue, producing cue-specific SI-SDRi, PESQ, and classification metrics.
-All cue: Each prompt includes all available single relative cues presented in the target speech of a mixture.
-The model is evaluated on mixtures with multiple cue types simultaneously, allowing us to assess performance under the combination of all available cues.
-Random cue: Each prompt contains a random subset of relative cues, and evaluation is performed on mixtures corresponding to these randomly selected cues.
-This measures the robustness of the two-stage method under partial or unpredictable cue guidance.
-The two-stage TSE also supports target speaker enrollment utterances as guidance, where the separated signal with the highest cosine similarity to the enrollment embedding is selected as the target.
-This enables a direct comparison with text-based relative cues.
-From the results, the TF-Locoformer-based single-stage TSE consistently outperforms the DPT-based single-stage TSE, achieving significant 3+ dB improvements in SI-SDRi and 0.7–1.1 improvements in PESQ, except for the age category, which remains challenging for both models.
-This suggests that age prediction performance is bottlenecked by the cue discriminability rather than the model architecture.
-The classification performance of different cue types reflects their relative contribution to the single-stage TSE baselines.
-Compared with the single-stage TF-Locoformer-based TSE, the two-stage method demonstrates clear advantages.
-It achieves over 3 dB SI-SDRi improvement for the random cue, all cue, language, transcription, gender, and speaking rate cues, and gains more than 2 dB SI-SDRi for the remaining cue types (except for the age category cue).
-Additionally, it provides an average improvement of over 0.5 in PESQ across all cue types, highlighting its strong practical potential.
+IV-C Two-Stage Method versus Single-Stage Baselines
+Experimental Setup: We first evaluate the proposed two-stage method against single-stage baselines.
+In the first stage, the separation model is trained from scratch to obtain separated speech signals.
+In the subsequent classification stage, a single model is trained and evaluated using template prompts for each mixture, with each prompt corresponding to one of the fourteen relative cue types listed in Table III.
+The single-stage baselines include a DPT-based TSE model from our prior work [7] and a TF-Locoformer-based TSE model, obtained by adapting the TF-Locoformer speech separation model for target speech extraction.
+Specifically, the model is reconfigured to predict only the target speaker by reducing the number of output speakers in the transposed convolution layer from 2 to 1.
+Text prompts categorized as “same” or “similar” are excluded from training, validation, and testing, as they provide limited discriminative power.
+We also train and evaluate a classification model to test an audio-based TSE approach.
+Here, the text encoder in the two-stage system is replaced with the audio encoder.
+For each mixture, a clean utterance from the same target speaker is randomly selected, ensuring that the same utterance is not always used for a given speaker.
+This setup evaluates the effectiveness of using an enrollment audio cue instead of textual relative cues for target speaker extraction.
+The clean utterance is normalized to match the mixture length via truncation or zero-padding: signals longer than the target duration are randomly cropped, while shorter signals are zero-padded with randomly distributed offsets.
+It is then convolved with a RIR randomly drawn from the RIR set described in Section II, producing the reverberant enrollment utterance.
+To ensure a fair comparison with textual cues, the same number of mixture samples is used for training and validation.
+Enrollment utterances are distinct from the corresponding mixtures, and evaluation is conducted on mixture sets of the same size as those used for each textual cue, enabling a direct comparison between audio- and text-based TSE approaches.
+Results and Analysis: We evaluate target speech extraction using scale-invariant signal-to-distortion ratio improvement (SI-SDRi) [22] and perceptual evaluation of speech quality (PESQ) [38].
+For the TF-Locoformer-based two-stage framework, we report classification accuracy on separated speech (Sep ACC) and overall TSE performance (SI-SDRi/PESQ), computed as the mean score between predicted and ground-truth target speech under different relative cue conditions.
+Results are summarized in Table IV.
+We evaluate three prompt types introduced in Section II (Individual, All, and Random) on their corresponding mixtures.
+The two-stage framework also supports target speaker enrollment, selecting the separated signal with the highest cosine similarity to the enrollment embedding, enabling direct comparison with text-based cues.
+TF-Locoformer-based single-stage TSE consistently outperforms the DPT-based counterpart, achieving over 3 dB SI-SDRi and 0.7–1.1 PESQ improvements, except for the age cue, which remains challenging.
+This suggests that age prediction performance is limited by cue discriminability rather than model architecture.
+Classification performance across cue types reflects their contribution to single-stage baselines.
+Compared to single-stage TF-Locoformer-based TSE, the two-stage method shows clear gains: over 3 dB SI-SDRi improvement for random, all, language, transcription, gender, and speaking rate cues, and over 2 dB for most others (except age).
+It also yields an average PESQ improvement exceeding 0.5, demonstrating strong practical potential.
+Some relative cues, such as gender, emotion, and pitch, perform competitively compared with enrollment audio.
+Interestingly, certain cue types—including all cues, random cues, transcription, loudness, distance, temporal order, and speaking duration—can even outperform the target speaker’s enrollment audio (see Table V, Sep ACC: textual cues vs.
+enrollment audio).
+This suggests that enrollment speech may contain confounding information, as it could be recorded in a different environment or exhibit speech characteristics differing from the target speech.
+These findings highlight the practicality of using textual descriptions of relative cues, which not only address issues such as missing enrollment speech and privacy concerns but can also improve extraction performance by leveraging discriminative cue types present in the mixture.
 Figure 3: Training and Validation Loss Comparison for TF-Locoformer-based single-stage TSE and TF-Locoformer-based Separation models over epochs.
 The loss is measured such that lower (more negative) is better.
-Interestingly, some relative cue types—including all cue, random cue, language, gender, loudness, distance, temporal order and speaking duration—can even outperform the enrollment speech of the target speaker.
-This suggests that enrollment speech may contain potentially confounding information, as it could be recorded in a different environment and exhibit speech characteristics that differ from the target speech.
-These findings further highlight the practicality of using textual descriptions of relative cues, which not only address issues such as missing enrollment speech and privacy concerns but can also improve extraction performance by leveraging some distinguishable cue types present in the mixture speech samples.
 We also present the training and validation loss curves for the TF-Locoformer-based single-stage TSE and TF-Locoformer-based separation models in Fig.
@@
 These observations highlight the advantage of the two-stage approach.
-To establish an oracle benchmark, we additionally train and evaluate a separate classification model on the target–inference reverberant–clean speech pairs from each mixture described in Section IV (Audio Mixtures of Dataset).
-The evaluation results are shown in Table VI.
-TABLE VI: Performance across different cues
-Cue Type Clean ACC Sep ACC
-Random Cue 99.3% 99.2%
-All Cue 99.8% 99.8%
-Language 99.5% 99.3%
-Transcription 90.8% 90.4%
-Gender 98.0% 98.1%
-Emotion 85.8% 87.2%
-Pitch Level 95.2% 94.6%
-Pitch Range 88.4% 88.7%
-Loudness 98.8% 97.2%
-Distance 99.0% 98.4%
-Age Category 64.0% 64.3%
-Temporal Order 100% 99.8%
-Speaking Rate 91.5% 91.3%
-Speaking Duration 96.3% 95.9%
-Enrollment Audio 96.1% 95.9%
-In general, classification performance on clean speech (Clean ACC) generally exceeds that on separated speech (Sep ACC), as expected.
-Interestingly, the accuracy for each cue type closely matches that of the model trained on separated speech generated by the TF-Locoformer separation model, demonstrating the high quality of the separated signals.
-The varying classification accuracies across cue types also provide insight into why some cues exert a stronger influence on target speech extraction, while others have a weaker effect.
-Figure 4: Overall trend of classification accuracy with respect to the attribute differences between target and interfering speakers for each cue type.
+To establish an oracle benchmark, we train separate classification models on reverberant–clean speech pairs (Section II, Audio Mixtures of Dataset) using both relative cues and enrollment utterances.
+We use identical prompt–sample pairs as in the separated-speech setting.
+The evaluation results are reported in Table V.
+TABLE V: Performance of classification across different cues
+Cue Type Textual Cues Enrollment Audio
+Clean ACC Sep ACC Clean ACC Sep ACC
+Random Cues 99.3% 99.2% 95.7% 95.7%
+All Cues 99.8% 99.8% 95.7% 95.7%
+Language 99.5% 99.3% 96.6% 96.6%
+Transcription 90.8% 90.4% 95.8% 95.9%
+Gender 98.0% 98.1% 98.2% 98.1%
+Emotion 85.8% 87.2% 88.4% 88.6%
+Pitch 95.2% 94.6% 95.8% 95.8%
+Pitch Range 88.4% 88.7% 95.9% 96.0%
+Loudness 98.8% 97.2% 95.7% 95.7%
+Distance 99.0% 98.4% 96.2% 96.4%
+Age 64.0% 64.3% 93.4% 93.4%
+Temporal Order 100% 99.8% 95.6% 95.6%
+Speaking Rate 91.5% 91.3% 95.1% 95.4%
+Speaking Duration 96.3% 95.9% 95.7% 95.6%
+Average 95.8% 95.6% 95.8% 95.8%
+Interestingly, the accuracy for each cue type (relative cues and enrollment audio) closely matches that of the model trained on speech separated by the TF-Locoformer model, demonstrating the high quality of the separated signals.
+The enrollment audio cue shows slightly more consistent classification performance between clean and separated speech, indicating that speaker identity information is more robust than that conveyed by some relative cues, as expected.
+The variation in classification accuracy across cue types also provides insight into why some cues exert a stronger influence on target speech extraction, while others have a weaker effect.
+Figure 4: classification accuracy as a function of the attribute difference between target and interfering speakers for each cue type.
 We further analyze classification accuracy as a function of the continuous-valued attribute difference between paired target–interference clean speech samples for each cue type.
@@
 4, accuracy exhibits a clear increasing trend as attribute differences grow.
-Except for age category, all cues reach near-perfect classification when the differences become sufficiently large.
+Except for age, all cues reach near-perfect classification when the differences become sufficiently large.
 IV-D Relative Cues versus Independent Cues
-Experimental Setup: We evaluate whether relative cues outperform independent cues for target speech extraction within a two-stage framework.
+Experimental Setup: We then evaluate whether relative cues outperform independent cues for target speech extraction within a two-stage framework.
 For discrete-valued attributes, relative and independent cues provide the same level of discriminability; therefore, we restrict our evaluation to originally continuous-valued attributes.
-In total, eight cues are considered, of which two—age category and temporal order—are excluded.
-The age category cue is excluded because the training data are limited to speakers aged 18–50, making absolute age categorization unreliable.
-Applying common-sense age groupings (young: <25<25, youth: 25–40, middle: 41–64, senior: >65>65) would lead to severe class imbalance (youth: 10,310; young: 7,063; middle: 2,627; senior: 0).
+In total, eight cues are considered, of which two—age and temporal order—are excluded.
+The age cue is excluded because the training data are limited to speakers aged 18–50, making absolute age categorization unreliable.
+Applying common-sense age groupings (young: <25<25, youth: 25–40, middle: 41–64, senior: >65>65) would lead to severe class imbalance (young: 7063; youth: 10310; middle: 2627; senior: 0).
 The temporal order cue, defined as the difference in onset time between the target and interfering speech, is inherently relational and therefore does not admit a meaningful absolute categorization.
-Consequently, the evaluation focuses on the remaining six cues: pitch level, pitch range, loudness, distance, speaking rate, and speaking duration.
-Separated speech samples are selected from the TF-Locoformer separation model corresponding to these cues.
+Consequently, the evaluation focuses on the remaining six cues: pitch, pitch range, loudness, distance, speaking rate, and speaking duration.
+Separated speech samples are selected from the TF-Locoformer separation model based on mixtures associated with these cue types.
 Two binary classification models are trained: one guided by relative-cue text prompts and the other by independent-cue text prompts.
-TABLE VII: Metrics for Independent and Relative Cues
+Independent Cues
+Independent cue categories are defined based on the distribution of attribute values in the training samples (including both target and interference speech), following previous studies[16, 17].
+We use equal-frequency binning, where each attribute is quantized into two or three categories with similar amount of samples per each category:
+  • Pitch: low / normal / high
+  • Pitch range: narrow / normal / wide
+  • Loudness: quiet / normal / loud
+  • Distance: near / far
+  • Speaking rate: slow / normal / fast
+  • Speaking duration: short / long
+Relative Cues
+Relative cue categories are defined based on the process explained in Section II and the thresholds specified in Table I:
+  • Pitch: lower / similar / higher
+  • Pitch range: narrower / similar / wider
+  • Loudness: quieter / similar / louder
+  • Distance: nearer / similar / farther
+  • Speaking rate: slower / similar / faster
+  • Speaking duration: shorter / similar / longer
+TABLE VI: Metrics for Independent and Relative Cues
 Independent Cues Relative Cues
@@
 Loudness all 10000 79.2% $6.1$ $2.94$ 10000 74.4% $3.7$ $2.85$
-Pitch Level non-same / non-similar 6630 98.1% $16.8$ $3.39$ 8909 95.4% $15.2$ $3.32$
-Pitch Level all 9990 82.4% $8.2$ $3.03$ 9990 90.4% $12.4$ $3.18$
+Pitch non-same / non-similar 6630 98.1% $16.8$ $3.39$ 8909 95.4% $15.2$ $3.32$
+Pitch all 9990 82.4% $8.2$ $3.03$ 9990 90.4% $12.4$ $3.18$
 Pitch Range non-same / non-similar 6652 89.7% $12.0$ $3.19$ 7612 89.6% $11.9$ $3.17$
@@
 Speaking Duration all 10000 64.2% 1.2-1.2 $2.61$ 10000 86.3% $10.4$ $3.10$
-Independent Cues
-Independent cue categories are defined based on the distribution of attribute values in the training samples (including both target and interference speech), following previous studies[17, 18].
-We use equal-frequency binning, where each attribute is quantized into two or three categories with similar amount of samples per each category:
-  • Pitch level: low / normal / high
-  • Pitch range: narrow / normal / wide
-  • Loudness: quiet / normal / loud
-  • Distance: near / far
-  • Speaking rate: slow / normal / fast
-  • Speaking duration: short / long
-Relative Cues
-Relative cue categories are defined by comparing the target speech with the interference speech:
-  • Pitch level: lower / similar / higher
-  • Pitch range: narrower / similar / wider
-  • Loudness: quieter / similar / louder
-  • Distance: nearer / similar / farther
-  • Speaking rate: slower / similar / faster
-  • Speaking duration: shorter / similar / longer
-We use the audio data corresponding to six relative cues—pitch level, pitch range, loudness, distance, speaking rate, and speaking duration—from the dataset described in Section IV.
-To isolate the effect of cue category differences, we generate prompts for independent cues and regenerate prompts for relative cues using predefined templates.
-Each prompt describes a single cue, and all prompts share the same sentence structure, differing only in whether the cue is expressed in relative or independent terms.
-For example, a relative-cue prompt is “Please extract the speaker with a higher pitch level in the audio,” whereas the corresponding independent-cue prompt is “Please extract the speaker with a high pitch level in the audio.”
-During training, samples in which both speech signals belong to the same absolute category (or the Similar relative category) are excluded.
-During evaluation, all samples are retained, and target labels are derived from the PIT results of the separation model.
+We use audio mixtures from the dataset described in Section IV that are annotated with the six relative cues—pitch, pitch range, loudness, distance, speaking rate, and speaking duration.
+We reuse the same predefined templates structure as described in Section II to generate prompts for independent cues.
+To isolate the effect of cue representation, we control for linguistic variability by regenerating relative-cue prompts using the same predefined templates as those used for independent cues.
+This ensures that both prompt types share identical sentence structure and phrasing, differing only in whether the cue is expressed in relative or independent terms.
+Consequently, any performance differences can be attributed specifically to the cue formulation rather than to variations in prompt wording or style.
+For example, a regenerated relative-cue prompt for a given target speech sample is “Please extract the speaker with a higher pitch in the audio,” whereas the corresponding independent-cue prompt is “Please extract the speaker with a high pitch in the audio.”
+During training, pairs of audio samples and prompts in which both speech signals belong to the same absolute category or the “similar” relative category are excluded.
+During evaluation, all pairs of audio samples and prompts are retained, with target labels derived from the PIT results of the separation model.
 Classification accuracy is reported under different conditions.
@@
 We further analyze cases in which target–interference pairs fall into different relative-cue categories but are quantized into the same, adjacent, or distinct absolute categories, highlighting scenarios where independent cues provide limited discriminability while relative cues remain informative.
+Results and Analysis: We compare overall classification accuracy across cue types to evaluate the effectiveness of relative versus independent cues for TSE.
+For independent cues, pairs in the same category are labeled “same,” and those in different categories “non-same.” For relative cues, pairs in the same category are labeled “similar,” and those in different categories “non-similar.” Results are reported over six continuous-valued attributes.
+For independent cues, we report classification accuracy along with SI-SDRi and PESQ for the “non-same” and “all” groups.
+For relative cues, we report these metrics for the “non-similar” and “similar” groups.
+We explicitly account for the number of samples (N) in each group to fairly compare whether relative cues can exploit more information than independent cues.
+Notably, each relative cue and its corresponding independent cue are evaluated using the same number of samples.
+The results are summarized in Table VI.
+The overall classification accuracy, SI-SDRi, and PESQ achieved with relative cues surpass those obtained with independent cues.
+Moreover, the number of samples in the non-similar group is larger than that in the similar group, highlighting the advantage of relative cues in leveraging more discriminative information.
+In particular, for pitch, pitch range, speaking rate, and speaking duration, the relative categories demonstrate a clear advantage over absolute categories in utilizing more distinguishing information.
 (a) Independent Speaking Rate Cue
@@
 Figure 5: Classification accuracy and corresponding sample distribution across regions for independent (a) and relative (b) speaking rate cues, where $x^{(\mathit{tar})}$ (y-axis) and $x^{(\mathit{inf})}$ (x-axis) represent the speaking rate values of the target and interfering speakers, respectively.
-Results and Analysis: Overall classification accuracy is compared across cue types to evaluate the effectiveness of relative versus independent cues for TSE.
-To facilitate analysis, pairs are grouped as follows: for independent cues, pairs in the same category are labeled the same group, and pairs in different categories are labeled the non-same group; for relative cues, pairs in the same category are labeled the similar group, and pairs in different categories are labeled the non-similar group.
-Results are presented across six continuous-valued attributes.
-For independent cues, we report the classification accuracy, as well as the resulting SI-SDRi and PESQ performance for the non-same and same groups.
-For relative cues, we report the classification accuracy and the corresponding SI-SDRi and PESQ performance for the non-similar and similar groups.
-We explicitly account for the number of samples (N) in each group to fairly compare whether relative cues can exploit more information than independent cues.
-Notably, each relative cue and its corresponding independent cue are evaluated using the same number of samples.
-The results are summarized in Table VII.
-The overall classification accuracy, SI-SDRi, and PESQ achieved with relative cues surpass those obtained with independent cues.
-Moreover, the number of samples in the non-similar group is larger than that in the similar group, highlighting the advantage of relative cues in leveraging more discriminative information.
-In particular, for pitch level, pitch range, speaking rate, and speaking duration, the relative categories demonstrate a clear advantage over absolute categories in utilizing more distinguishing information.
-For the distance and loudness cues, using relative categories results in more samples being assigned to the similar group than when using absolute categories.
-This is due to the limited ranges of the simulated speaker-to-microphone distances and SIR values.
-In our dataset, the speaker-to-microphone distances in the simulated mixtures range from 0.3 to 1.5 m, and an auditory distance threshold of 0.5 m is used to distinguish between the two speakers.
-Consequently, many target-interference speech pairs are classified as having similar distances.
+For the distance and loudness cues, using relative categories assigns more samples to the “similar” group than absolute categories.
+This is due to the limited ranges of simulated speaker-to-microphone distances and SIR values.
+In our dataset, distances range from 0.3 to 1.5 m, with a 0.5 m threshold for distinguishing speakers, leading many pairs to be classified as having similar distances.
 A similar effect is observed for the loudness cue: loudness is approximated using SIR, which is constrained to the range of -6 to 6 dB in our simulation.
 With a loudness cue threshold of 3 dB, nearly half of the samples are categorized as having similar loudness.
+It is informative to examine the specific categories of target–interference samples.
+Fig.
+5 presents the classification accuracy and sample distribution for both independent (Fig.
+5(a)) and relative (Fig.
+5(b)) cues, using speaking rate as an example.
+Compared with the relative speaking rate cue, the independent cue results in more samples being categorized into the “same” group (slow–slow, normal–normal, fast–fast), which cannot be distinguished.
+In contrast, the relative cue preserves the fine-grained distinctions among these samples, supporting the assumption in Section II.
+For the independent speaking rate cue, accuracy is highest when target and interference fall into clearly distinct categories (e.g., slow–fast or fast–slow), and decreases for adjacent categories (e.g., slow–normal or normal–fast), as smaller relative differences make classification more difficult.
 Figure 6: Comparison between independent and relative cue categorizations for the speaking rate attribute.
@@
 The gray region denotes target–interference speaking-rate differences within $\pm 15\%$.
-It is informative to examine the specific categories of target–interference samples.
-Fig.
-5 presents the classification accuracy and sample distribution for both relative and independent cues, using speaking rate as an example.
-When the target and interference speech fall into clearly distinct categories (e.g., slow–fast or fast–slow), the model achieves the highest accuracy.
-In contrast, accuracy decreases when they belong to adjacent categories (e.g., slow–normal or normal–fast), supporting the assumption in Section II.
-To further analyze the relationship between independent and relative cue categorizations, we focus on target–interference pairs labeled as relative-different and examine whether they fall into the same, adjacent, or distinct absolute categories.
+To further analyze the relationship between independent and relative cue categorizations, we focus on target–interference pairs labeled as “non-similar” and examine whether they fall into the same, adjacent, or distinct absolute categories.
 Using speaking rate as an example (Fig.
-6), we report classification accuracy under these three absolute conditions for both the full set and the relative-different subset, along with their sample distribution.
-Accuracy increases monotonically with the magnitude of speaking rate difference between two speakers.
-In the same condition, performance is relatively low (48.9%), but restricting to relative-different pairs improves accuracy to 79.4%, indicating that relative comparison provides additional discriminative information even when absolute categories are identical.
-In the adjacent condition, accuracy is high (87.3%) and slightly improves to 89.5% under the relative-different subset.
-In the distinct condition, performance reaches 98.3% in both cases.
-These results show that classification performance strongly correlates with speaking rate difference between two speakers, and that relative cues particularly benefit challenging cases where absolute quantization obscures pairwise distinctions.
+6), we report classification accuracy for these conditions on both the full set and the relative-different subset, along with their sample distribution.
+Accuracy increases with speaking rate difference: in the same condition, it is 48.9% but improves to 79.4% for “non-similar” pairs, indicating added discriminative power from relative comparison.
+In the adjacent condition, accuracy rises from 87.3% to 89.5%, while in the distinct condition it reaches 98.3% in both cases.
+These results show that performance strongly correlates with speaking rate difference, with relative cues particularly benefiting cases where absolute quantization obscures pairwise distinctions.
 Fig.
@@
 We have presented a comprehensive study on text-guided target speech extraction that leverages inter-speaker relative cues across a wide range of speech attributes.
-By providing a theoretical analysis, redesigning the model architecture with a two-stage separation-and-classification approach, and empirically evaluating relative cues versus independent cues on comparable continue-valued attributes.
-Experimental results show that the proposed two-stage approach substantially improves both signal-level (SI-SDRi) and perceptual (PESQ) performance, outperforming single-stage text-based TSE.
-Analysis of classification accuracy across cue types indicates that while some cues have strong discriminative power, certain text cues are even more effective than conventional audio-based methods, highlighting their practical potential.
-Finally, our results confirm that relative cues generally outperform independent cues, underscoring their advantage for text-guided target speech extraction.
-Notably, adapting a state-of-the-art separation model into a single-stage TSE system can also significantly enhance target speech extraction, further validating the effectiveness of leveraging advanced separation models in TSE frameworks
-Overall, our findings underscore the effectiveness of integrating perceptually relative cues with a two-stage model for robust and flexible text-guided TSE.
-The proposed framework not only mitigates speaker confusion inherent in single-stage systems but also provides valuable insights into the relative importance of different speech attributes for TSE.
+This is achieved through theoretical analysis, redesigning the model architecture with a two-stage separation-and-classification approach, and empirically evaluating relative versus independent cues on comparable continuous-valued attributes.
+Experimental results demonstrate that the proposed two-stage approach substantially improves both signal-level (SI-SDRi) and perceptual (PESQ) performance, outperforming single-stage text-based TSE.
+Analysis of classification accuracy across cue types indicates that, while some cues have strong discriminative power, certain textual cues are even more effective than conventional audio-based methods, highlighting their practical potential.
+Furthermore, our results confirm that relative cues generally outperform independent cues, underscoring their advantage for text-guided target speech extraction.
+Notably, adapting a state-of-the-art separation model into a single-stage TSE system can also significantly enhance target speech extraction, further validating the effectiveness of incorporating advanced separation models into TSE frameworks.
+Overall, our findings highlight the effectiveness of integrating perceptually relative cues with a two-stage model for robust and flexible text-guided TSE.
+The proposed two-stage framework not only mitigates the speaker confusion inherent in single-stage systems but also provides valuable insights into the relative importance of different speech attributes for TSE.